### v1

# non-asymptotic estimates for markov transition matrices with rigorous error bounds

Abstract.
We establish non-asymptotic error bounds for the classical Maximal Likelihood Estimation of the transition matrix of a given Markov chain. Meanwhile, in the reversible case, we propose a new reversibility-preserving online Symmetric Counting Estimation of the transition matrix with non-asymptotic deviation bounds. Our analysis is based on a convergence study of certain Markov chains on the length-2 path spaces induced by the original Markov chain.
1School of Mathematical Sciences, Peking University. E-mail: dhuang@math.pku.edu.cn
2School of Mathematical Sciences, Peking University. E-mail: lixiangyuan23@stu.pku.edu.cn
1. introduction
Markov models are fundamental in various fields for their efficiency in modeling time-dependent scenarios where the future state depends solely on the present state. The inverse problem of estimating the transition matrix of a given Markov chain is one of the most fundamental tasks in many scientific applications [15, 5, 9, 18, 14].
Let $\Omega$ be a finite state space, and let $\bm{P}=[p(u,v)]_{u,v\in\Omega}$ be an irreducible transition matrix with an invariant distribution $\mu$. Our goal is to estimate the transition probability $p(u,v)$ for all pairs of states when $\bm{P}$ is unknown. Suppose we have access to a finite Markov sequence $\{u_{k}\}_{k=1}^{n+1}$ generated by $\bm{P}$. An extensively studied method, known as the Maximal Likelihood Estimation (MLE), simply counts the number of jumps from $u$ to $v$ for all $u,v\in\Omega$ and uses the empirical mean of the indicator function $\mathbb{1}_{(u,v)}(\cdot,\cdot)$ to approximate $\mu(u)p(u,v)$:
$r_{n}(u,v):=\frac{1}{n}\sum\limits_{k=1}^{n}\mathbb{1}_{(u,v)}\left(u_{k},u_{k+1}\right)\approx\mu(u)p(u,v).$ (1.1)
Then $\mu(u)$ and $p(u,v)$ can be approximated by $\sum_{w\in\Omega}r_{n}(u,w)$ and $r_{n}(u,v)/\sum_{w\in\Omega}r_{n}(u,w)$, respectively. We will use $\bm{R}_{n}=[r_{n}(u,v)]_{u,v\in\Omega}$ to denote the matrix form of MLE. It is shown in [1, 3] that the error of MLE is asymptotically $\mathcal{O}(1/\sqrt{n})$ by the central limiting theorem for Markov chains. However, to the best of our knowledge, there is no quantitative result on the non-asymptotic behavior of MLE in the current literature.
Our first contribution is to establish non-asymptotic error bounds of MLE. To achieve this, we interpret MLE as the Markov empirical mean of the indicator function $\mathbb{1}_{(u,v)}(\cdot,\cdot)$ corresponding to some transition matrix $\bm{P}_{2}$ defined on the length-2 path space $\Omega^{2}$. We then study the spectrum of $\bm{P}_{2}$ and apply concentration inequalities for Markov chains to prove the following entry-wise tail bound for the error of MLE.
Theorem 1.1.
For any $t>0$ and any $u,v\in\Omega$,
$\mathbb{P}\left\{|r_{n}(u,v)-\mu(u)p(u,v)|\geq t\right\}\leq C_{1}\|\nu/\mu\|_{\infty}\exp\left(-C_{2}\eta_{s}(\bm{P})nt^{2}\right),$ (1.2)
where $C_{1}$, $C_{2}$ are absolute constants, $\nu$ is the initial distribution of $u_{1}$, and $\eta_{s}(\bm{P})$ is the symmetric spectral gap of $\bm{P}$.
There has been extensive research on concentration/deviation inequalities for Markov chains [11, 12, 13, 17, 6, 10]. Most of these results presented in these studies typically require the transition matrix to satisfy certain spectral conditions, which $\bm{P}_{2}$ fails to meet. We also notice that the proofs of concentration inequalities for general Markov chains given in [6, 10] seem to have some gap. Therefore, we choose to adopt the result in [8], that can be directly applied to $\bm{P}_{2}$.
In numerous applications, the Markov chain considered is reversible [7], i.e. the transition matrix satisfies the detailed balance condition. In this case, it is important that the estimator preserves the reversibility of $\bm{P}$  [2, 18, 14]. Similarly, the theories proposed in the literature do not provide any non-asymptotic convergence rate guarantee for those estimators.
Our second contribution is to propose an online reversibility-preserving Symmetric Counting Estimation (SCE) method as follows in the reversible case. We construct a reversible Markov chain $\{\tilde{\bm{u}}_{k}\}_{k\geq 1}$ on the symmetrized length-2 path space $\tilde{\Omega}^{2}$ and use its empirical mean under some matrix-valued function $\tilde{\bm{F}}$ to approximate $\bm{D}_{\mu}\bm{P}$:
$\bm{H}_{n}:=\frac{1}{n}\sum\limits_{k=1}^{n}\tilde{\bm{F}}(\tilde{\bm{u}}_{k})\approx\bm{D}_{\mu}\bm{P}.$ (1.3)
Let $\tilde{\bm{P}}_{2}$ denote the corresponding transition matrix on $\tilde{\Omega}^{2}$. A non-asymptotic error bound of SCE in the matrix form can be obtained by exploiting the spectral properties of $\tilde{\bm{P}}_{2}$ and applying concentration inequalities for Markov dependent random matrices recently established in [16].
Theorem 1.2.
For any $t>0$,
$\mathbb{P}\left\{\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|\geq t\right\}\leq C_{1}d^{2-\pi/4}\|\nu/\mu\|_{\infty}\exp\left(-C_{2}\eta(\bm{P})nt^{2}\right),$ (1.4)
where $d=|\Omega|$, $C_{1}$, $C_{2}$ are absolute constants, $\nu$ is some initial distribution on $\Omega$, and $\eta(\bm{P})$ is the spectral gap of $\bm{P}$.
Non-asymptotic expected error bounds are sometimes more favorable in applications. A straightforward way to bound the error expectation is by integrating the corresponding tail bound. However, for MLE and SCE, the expectation bounds derived in this way both depend on the dimension $d$, which is inconsistent with our numerical observations. Instead, we derive dimension-free expectation bounds of both estimators by utilizing the special structure $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$.
Theorem 1.3.
It holds that
$\operatorname{\mathbb{E}}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}\right]=\mathcal{O}\left(\frac{1}{\sqrt{n\eta_{s}(\bm{P})}}\right),$ (1.5)
where $\eta_{s}(\bm{P})$ is the symmetric spectral gap of $\bm{P}$. Moreover, when $\bm{P}$ is reversible,
$\operatorname{\mathbb{E}}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}\right]=\mathcal{O}\left(\frac{1}{\sqrt{n\eta(\bm{P})}}\right),\quad\operatorname{\mathbb{E}}\left[\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}\right]=\mathcal{O}\left(\frac{1}{\sqrt{n\eta(\bm{P})}}\right).$ (1.6)
Our numerical results suggest that for each of the two methods, the mean square error (MSE) is independent of $d$, and that 1/MSE depends linearly on $n$ and the spectral gap, which justifies the optimality of our theoretical results.
The rest of this paper is organized as follows. In Section 2, we develop some basic facts about Markov chains and state our main results which are complete versions of the theorems above. Section 3 focuses on constructing the induced Markov chains on path spaces and exploring their properties. We then use these properties to prove our main results in Section 4. The presentation of numerical results in Section 5 concludes the paper.
2. non-asymptotic estimates of markov matrices
In this section, we outline the main results of this paper with proofs appearing in Section 4, before which we develop some basic facts about Markov chains.
2.1. Fundamentals
2.1.1. Notation
Throughout this paper, matrices are written in uppercase bold letters. In particular, $\bm{E}_{uv}$ denotes the matrix with 1 at position $(u,v)$ and 00 elsewhere, $\bm{\mathbf{I}}_{d}$ denotes the $d$-dimensional identity matrix, and $\bm{1}_{d}$ denotes the d-dimensional all-one vector. In cases where it does not cause ambiguity, we will omit the subscript $d$ to use $\bm{\mathbf{I}}$ and $\bm{1}$ instead of $\bm{\mathbf{I}}_{d}$ and $\bm{1}_{d}$.
For a real vector $\alpha\in\mathbb{R}^{d}$, we use $\|\alpha\|_{\infty}$ for the $L_{\infty}$ norm of $\alpha$ and $\bm{D}_{\alpha}$ for the $d\times d$ diagonal matrix with the elements of $\alpha$ on the diagonal. For a matrix $\bm{A}\in\mathbb{R}^{m\times n}$, we write $\|\bm{A}\|$ for the $l_{2}$ operator norm, $\|\bm{A}\|_{\mathrm{F}}$ for the Frobenius norm, $\bm{A}(u,v)$ for the entry at position $(u,v)$, and $\bm{A}^{\mathrm{T}}$ for the transpose. For $\bm{A}\in\mathbb{R}^{d\times d}$, we write $\operatorname{spec}(\bm{A})$ for the spectrum set and $\operatorname{tr}[\bm{A}]$ for the trace.
Let $\mathbb{H}_{d}$ denote the linear space of all $d\times d$ real symmetric matrices and $\mathbb{H}_{d}^{+}$ denote the cone of all positive-semidefinite matrices. We use the symbol $\preceq$ for the semidefinite partial order on Hermitian matrices: for matrices $\bm{A},\bm{B}\in\mathbb{H}_{d}$, the inequality $\bm{A}\preceq\bm{B}$ means that $\bm{B}-\bm{A}\in\mathbb{H}_{d}^{+}$.
Let $\Omega$ be a Polish space equipped with a probability measure $\mu$. Define $\mathbb{E}_{\mu}$ the expectation with respect to the measure $\mu$. When applied to a random matrix, $\mathbb{E}_{\mu}$ computes the entry-wise expectation. In this paper, the number of states $|\Omega|$ is assumed to be finite and is always denoted by $d$. Any real-valued function on $\Omega$ can be seen as a finite-dimensional real vector and vice versa.
2.1.2. Foundation of Markov chains
Let $\{u_{k}\}_{k\geq 1}$ be a discrete-time Markov chain on a finite state space $\Omega$, and let $\bm{P}=[p(u,v)]_{u,v\in\Omega}$ denote the corresponding transition matrix. Let $\mu$ be an invariant distribution of $\bm{P}$ in the sense that
$\sum\limits_{u\in\Omega}\mu(u)p(u,v)=\mu(v),\quad\text{ for all}\,\ v\in\Omega.$ (2.1)
For a function $f:\Omega\rightarrow\mathbb{R}$, the action of $\bm{P}$ as a linear operator on $f$ is defined as
$\bm{P}f(u)=\mathbb{E}\left[f(u_{2})|u_{1}=u\right]=\sum\limits_{v\in\Omega}p(u,v)f(u).$ (2.2)
We say $\{u_{k}\}_{k\geq 1}$ is reversible with respect to $\mu$ if
$\mathbb{E}_{\mu}\left[g\cdot\bm{P}f\right]=\mathbb{E}_{\mu}\left[f\cdot\bm{P}g\right],\quad\text{for any}\,\ f,g:\Omega\rightarrow\mathbb{R}.$ (2.3)
The reversibility of $\{u_{k}\}_{k\geq 1}$ is equivalent to the detailed balance condition of $\bm{P}$:
$\mu(u)p(u,v)=\mu(v)p(v,u),\quad\text{for any}\,\ u,v\in\Omega.$ (2.4)
In this case, we also say that $\bm{P}$ is reversible.
2.1.3. Absolute spectral gap and symmetric spectral gap
Throughout this paper, we will assume that the transition matrix $\bm{P}$ is irreducible, and thus its invariant distribution $\mu$ is unique. Let $L_{2,\mu}$ be the Hilbert space of all real-valued functions on $\Omega$ endowed with the inner product
$\langle f,g\rangle_{\mu}=\mathbb{E}_{\mu}\left[f\cdot g\right],$ (2.5)
and $L_{2,\mu}^{0}=\{f\in L_{2,\mu}:\operatorname{\mathbb{E}}_{\mu}[f]=0\}$ be the mean-zero subspace of $L_{2,\mu}$. We will use $\|\cdot\|_{\mu}$ to denote both the $L_{2,\mu}$ norm of a vector and the $L_{2,\mu}$ operator norm of a matrix. Write $\bm{\Pi}=\bm{1}\mu^{\mathrm{T}}$, the projection onto $\bm{1}$ under the $L_{2,\mu}$ inner product. One can verify that $\bm{\Pi}\bm{P}=\bm{P}\bm{\Pi}=\bm{\Pi}$, and therefore
$\bm{P}^{k}-\bm{\Pi}=(\bm{P}-\bm{\Pi})^{k}=(\bm{P}-\bm{\Pi})^{k}(\bm{\mathbf{I}}-\bm{\Pi}),\quad\text{for any}\,\ k\in\mathbb{N}_{+}.$ (2.6)
We say the transition matrix $\bm{P}$ admits a nonzero absolute spectral gap $\eta_{a}(\bm{P})=1-\lambda_{a}(\bm{P})$ if
$\lambda_{a}(\bm{P}):=\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\|\bm{P}h\|_{\mu}}{\|h\|_{\mu}}=\|\bm{P}-\bm{\Pi}\|_{\mu}<1.$ (2.7)
Let $\bm{P}^{*}$ be the adjoint of $\bm{P}$ with respect to the $L_{2,\mu}$ inner product and $\bm{A}:=(\bm{P}+\bm{P}^{*})/2$ be the additive symmetrization of $\bm{P}$. We say $\bm{P}$ admits a nonzero symmetric spectral gap $\eta_{s}(\bm{P})=1-\lambda_{s}(\bm{P})$ if
$\lambda_{s}(\bm{P}):=\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\left\langle h,\bm{P}h\right\rangle_{\mu}}{\|h\|_{\mu}^{2}}=\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\left\langle h,\bm{A}h\right\rangle_{\mu}}{\|h\|_{\mu}^{2}}<1.$ (2.8)
By the irreducibility of $\bm{P}$ and the Perron-Frobenius theorem, we know $\mu(u)\neq 0$ for any $u\in\Omega$, and thus $\bm{D}_{\mu}$ is invertible. It is not hard to find that $\lambda_{a}(\bm{P})$ is the second largest singular value of the matrix $\bm{D}_{\mu}^{1/2}\bm{P}\bm{D}_{\mu}^{-1/2}$, which is also the largest singular value of $\bm{D}_{\mu}^{1/2}(\bm{P}-\bm{\Pi})\bm{D}_{\mu}^{-1/2}$, and $\lambda_{s}(\bm{P})$ is the second largest eigenvalue of $(\bm{D}_{\mu}^{1/2}\bm{P}\bm{D}_{\mu}^{-1/2}+\bm{D}_{\mu}^{-1/2}\bm{P}^{\mathrm{T}}\bm{D}_{\mu}^{1/2})/2$. By Jensen’s inequality,
$\|\bm{P}h\|_{\mu}^{2}=\sum\limits_{u\in\Omega}\mu(u)\left(\sum\limits_{v\in\Omega}p(u,v)h(v)\right)^{2}\leq\sum\limits_{u\in\Omega}\mu(u)\sum\limits_{v\in\Omega}p(u,v)h(v)^{2}=\|h\|_{\mu}^{2}.$ (2.9)
Hence $\lambda_{a}(\bm{P})\leq 1$. Moreover, a direct application of the Cauchy-Schwarz inequality implies $|\lambda_{s}(\bm{P})|\leq\lambda_{a}(\bm{P})$, and therefore $\eta_{s}(\bm{P})\geq\eta_{a}(\bm{P})$. One should realize that $\bm{A}$ is also a transition matrix (the sum of each row is $1$), and it is irreducible since $\bm{P}$ is irreducible. Therefore, we always have $\lambda_{s}(\bm{P})<1$. However, irreducibility does not guarantee that $\lambda_{a}(\bm{P})<1$. In fact, in Section 3 we will construct a family of irreducible transition matrices whose absolute spectral gaps are zero.
In particular, when $\bm{P}$ is reversible, $\lambda_{s}(\bm{P})$ is the second largest eigenvalue of $\bm{P}$ and $\eta_{s}(\bm{P})=1-\lambda_{s}(\bm{P})$ is the gap between the largest two eigenvalues of $\bm{P}$. In this case we will omit the subscript $s$ to use $\lambda(\bm{P})$ for the second largest eigenvalue of $\bm{P}$ and $\eta(\bm{P})=1-\lambda(\bm{P})$ for the spectral gap of $\bm{P}$.
For technical convenience, in the rest of this paper we will assume $\bm{P}$ is a dense matrix, i.e. $p(u,v)\neq 0$ for any $u,v\in\Omega$. Note that all spectral properties discussed above are well-posed under small perturbation on $\bm{P}$, and hence all arguments and conclusions in this paper can be readily generalized to non-dense cases by the limiting argument.
2.2. Main results
In this subsection, we outline the main results of this paper, which are complete versions of Theorem 1.1, 1.2 and 1.3 in the introduction. Let $\{u_{k}\}_{k\geq 1}$ be a Markov chain on $\Omega$ and $\bm{P}$ be the corresponding irreducible transition matrix we aim to estimate. The length-2 path space $\Omega^{2}$ is defined to be the Cartesian product of $\Omega$ with itself. The first contribution of the paper is to construct a Markov chain $\{\bm{u}_{k}\}_{k\geq 1}$ on $\Omega^{2}$ induced by $\{u_{k}\}_{k\geq 1}$ (see Section 3) and establish non-asymptotic error bounds of MLE by interpreting it as the empirical mean under some matrix-valued function $\bm{F}$ defined on $\Omega^{2}$:
$\bm{R}_{n}=\frac{1}{n}\sum\limits_{k=1}^{n}\bm{F}(\bm{u}_{k}).$ (2.10)
To be specific, in the following theorem we give entry-wise tail bounds and dimension-free expectation bounds for the error of MLE in terms of $n$ and spectral gap of $\bm{P}$.
Theorem 2.1.
Let $\{u_{k}\}_{k\geq 1}$ be a Markov chain on $\Omega$ generated by an irreducible transition matrix $\bm{P}$ with an invariant distribution $\mu$, and let $\bm{R}_{n}=[r_{n}(u,v)]_{u,v\in\Omega}$ be the matrix form of the corresponding MLE.
  1. (1)
    $\bm{R}_{n}$ converges to $\bm{D}_{\mu}\bm{P}$ almost surely.
  2. (2)
    Let $\nu$ be the initial distribution of $u_{1}$. For any $t>0$ and any $u,v\in\Omega$,
    $\mathbb{P}\left\{\left|r_{n}(u,v)-\mu(u)p(u,v)\right|\geq t\right\}\leq 2\|\nu/\mu\|_{\infty}\exp\left(\frac{-n\eta_{s}(\bm{P})t^{2}}{8\sqrt{\left(2+5\eta_{s}(\bm{P})/2\sqrt{2}\right)^{2}+t^{2}}}\right).$ (2.11)
  3. (3)
    It holds that
    $\operatorname{\mathbb{E}}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\leq\|\nu/\mu\|_{\infty}\min\left\{\frac{4+\eta_{s}(\bm{P})}{n\eta_{s}(\bm{P})},\frac{2+\eta_{a}(\bm{P})}{n\eta_{a}(\bm{P})}\right\}.$ (2.12)
    Moreover, when $\bm{P}$ is reversible,
    $\operatorname{\mathbb{E}}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\leq\|\nu/\mu\|_{\infty}\frac{2+\eta(\bm{P})}{n\eta(\bm{P})}.$ (2.13)
When $\bm{P}$ is reversible, the Markov chain constructed in Theorem 2.1 is nonreversible in general. In order to utilize the reversibility of $\bm{P}$, we can symmetrize $\Omega^{2}$ by viewing $(u_{1},u_{2})$ and $(u_{2},u_{1})$ as the same element to define a symmetric length-2 path space $\tilde{\Omega}^{2}$ and construct a reversible transition matrix $\tilde{\bm{P}}_{2}$ on $\tilde{\Omega}^{2}$. When we have access to an online oracle of $\bm{P}$ that takes an input $u\in\Omega$ and outputs a state $v\in\Omega$ with probability $p(u,v)$, we propose a Symmetric Counting Estimation(SCE) algorithm that generates a sequence $\{\tilde{\bm{u}}_{k}\}_{k\geq 1}\subset\tilde{\Omega}^{2}$ according to $\tilde{\bm{P}}_{2}$ and approximates $\bm{D}_{\mu}\bm{P}$ using the symmetric matrix empirical mean
$\bm{H}_{n}=\frac{1}{n}\sum\limits_{k=1}^{n}\tilde{\bm{F}}(\tilde{\bm{u}}_{k}),$ (2.14)
where $\tilde{\bm{F}}$ is some matrix-valued function defined on $\tilde{\Omega}^{2}$ to be specified later. We will show that $\tilde{\bm{P}}_{2}$ admits a nonzero absolute spectral gap that is exactly half of the spectral gap of $\bm{P}$. As a result, we can obtain non-asymptotic error bounds for SCE as in the following theorem.
Theorem 2.2.
Let $\bm{P}$ be an irreducible reversible transition matrix on $\Omega$ with an invariant distribution $\mu$, and let the corresponding symmetric path space $\tilde{\Omega}^{2}$ be defined as above. The followings hold:
  1. (1)
    There exists a reversible Markov chain $\{\tilde{\bm{u}}_{k}\}_{k\geq 1}$ on $\tilde{\Omega}^{2}$ and a symmetric matrix $\bm{H}_{n}=[h_{n}(u,v)]_{u,v\in\Omega}$ generated by $\{\tilde{\bm{u}}_{k}\}_{k\geq 1}$ such that $\bm{H}_{n}$ converges to $\bm{D}_{\mu}\bm{P}$ almost surely.
  2. (2)
    For any $t>0$,
    $\mathbb{P}\left\{\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|\geq t\right\}\leq 2\|\nu/\mu\|_{\infty}d^{2-\frac{\pi}{4}}\exp\left(\frac{-n\eta(\bm{P})t^{2}}{128\pi^{-2}(2-\eta(\bm{P})/2)+1024\pi^{-3}t}\right).$ (2.15)
    where $\nu$ is some initial distribution on $\Omega$ which is related to the generation of $\tilde{\bm{u}}_{1}$.
  3. (3)
    It holds that
    $\operatorname{\mathbb{E}}\left[\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\leq\|\nu/\mu\|_{\infty}\frac{4-\eta(\bm{P})}{n\eta(\bm{P})}.$ (2.16)
3. Markov chains on path spaces
This section focuses on the induced Markov chains on the path spaces. First, we construct Markov chains on $\Omega^{2}$ and $\tilde{\Omega}^{2}$ with $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$ as their transition matrices respectively. Following this, the framework of MLE and SCE is proposed in detail. Lastly, we study the spectral properties of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$, which are crucial for establishing non-asymptotic error bounds for MLE and SCE in Section 4.
3.1. Construction of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$
Let us introduce the construction of the induced Markov chains on length-2 path spaces.
3.1.1. Construction of $\bm{P}_{2}$
Recall that $\bm{P}=[p(u,v)]_{u,v\in\Omega}$ is a transition matrix on the state space $\Omega$ with an unique invariant distribution $\mu$ Then the length-2 path space $\Omega^{2}=\{(u_{1},u_{2}):u_{1},u_{2}\in\Omega\}$ is naturally equipped with a distribution $\mu_{2}$ that is given by
$\mu_{2}\left(u_{1},u_{2}\right)=\mu\left(u_{1}\right)p\left(u_{1},u_{2}\right),\quad u_{i}\in\Omega,i=1,2.$ (3.1)
We will use bold letters, e,g. $\bm{u}=(u_{1},u_{2})$ to denote elements in $\Omega^{2}$. This Markov transition law on $\Omega$ naturally induces a Markov transition law on $\Omega^{2}$, whose transition probabilities are as follows:
  1. (1)
    Given a state $\bm{u}=(u_{1},u_{2})\in\Omega$, delete the first element $u_{1}$.
  2. (2)
    Choose a state $v\in\Omega$ according to the probability
    $\mathbb{P}\{\text{choose}~{}v\}=p\left(u_{2},v\right),$ (3.2)
    and then jump to the state $\bm{v}=(u_{2},v)$.
We will use $\bm{P}_{2}=[p_{2}(\bm{u},\bm{v})]_{\bm{u},\bm{v}\in\Omega^{2}}$ to denote the transition matrix of this Markov chain on $\Omega^{2}$. Later we will show that $\mu_{2}$ is the unique invariant distribution of $\bm{P}_{2}$.
In general, suppose $\{u_{k}\}_{k\geq 1}$ is a Markov sequence generated by $\bm{P}$ with an initial distribution $u_{1}\sim\nu$. Then it is not hard to check that $\{\bm{u}_{k}=(u_{k},u_{k+1})\}_{k\geq 1}$ can be viewed as a Markov chain on $\Omega^{2}$ generated by $\bm{P}_{2}$ with an initial distribution $\bm{u}_{1}=(u_{1},u_{2})\sim\nu_{2}$, where
$\nu_{2}(u_{1},u_{2})=\nu(u_{1})p(u_{1},u_{2}).$ (3.3)
3.1.2. Construction of $\tilde{\bm{P}}_{2}$
Note that $\bm{P}_{2}$ is non-reversible in general even if $\bm{P}$ is reversible. Next we turn to the construction of a reversible Markov chain on the symmetric length-2 path space for a reversible $\bm{P}$. For two paths $\bm{u}=(u_{1},u_{2})$ and $\bm{v}=(v_{1},v_{2})\in\Omega^{2}$, we say $\bm{u}$ is equivalent to $\bm{v}$, denoted by $\bm{u}\sim\bm{v}$, if
$\left(u_{1},u_{2}\right)=\left(v_{1},v_{2}\right)\quad\text{or}\quad\left(u_{1},u_{2}\right)=\left(v_{2},v_{1}\right).$ (3.4)
We then define the symmetric path space as $\tilde{\Omega}^{2}:=\Omega^{2}/\sim$. For a path $\bm{u}\in\Omega^{2}$ and an equivalent class $\tilde{\bm{u}}\in\tilde{\Omega}^{2}$, we write $\bm{u}\sim\tilde{\bm{u}}$ or $\tilde{\bm{u}}\sim\bm{u}$ if $\bm{u}$ falls in $\tilde{\bm{u}}$. The distribution $\mu_{2}$ on $\Omega^{2}$ induces a distribution $\tilde{\mu}_{2}$ on $\tilde{\Omega}^{2}$:
$\tilde{\mu}_{2}(\tilde{\bm{u}}):=\sum\limits_{\bm{u}\sim\tilde{\bm{u}}}\mu_{2}(\bm{u}),\quad\tilde{\bm{u}}\in\tilde{\Omega}^{2}.$ (3.5)
Correspondingly, we can construct a reversible Markov chain on the symmetric path space $\tilde{\Omega}^{2}$ as follows whose invariant distribution is exactly $\tilde{\mu}_{2}$:
  1. (1)
    Given a state $\tilde{\bm{u}}=\widetilde{(u_{1},u_{2})}\in\tilde{\Omega}^{2}$, randomly choose a path $\bm{w}=(w_{1},w_{2})\in\Omega^{2}$ according to the probability
    $\mathbb{P}\{w_{1}=u_{1},w_{2}=u_{2}\}=\mathbb{P}\{w_{1}=u_{2},w_{2}=u_{1}\}=\frac{1}{2}.$ (3.6)
  2. (2)
    Choose a state $v\in\Omega$ according to the probability
    {choose $\mathbb{P}\{\text{choose }v\}=p\left(w_{2},v\right),$ (3.7)
    and then jump to the state $\tilde{\bm{v}}\in\tilde{\Omega}^{2}$ such that $(w_{2},v)\sim\tilde{\bm{v}}$.
We will use $\tilde{\bm{P}}_{2}=[\tilde{p}_{2}(\tilde{\bm{u}},\tilde{\bm{v}})]_{\tilde{\bm{u}},\tilde{\bm{v}}\in\tilde{\Omega}^{2}}$ to denote the transition matrix of this Markov chain on $\tilde{\Omega}^{2}$. Roughly speaking, $\tilde{\bm{P}}_{2}$ can be seen as a modification of $\bm{P}_{2}$ by randomly reversing the path with probability $1/2$, which makes the process reversible. Similarly, we will show that $\tilde{\mu}_{2}$ is the unique invariant distribution of $\tilde{\bm{P}}_{2}$.
Previously, we can get a Markov sequence of $\bm{P}_{2}$ directly from a Markov sequence of $\bm{P}$. But this is not the case for $\tilde{\bm{P}}_{2}$ due to the extra reversing operation. Nevertheless, from the construction of $\tilde{\bm{P}}_{2}$ we can see that, as long as we have access to an online oracle of $\bm{P}$ that takes in a state $u\in\Omega$ and outputs $v\in\Omega$ with probability $p(u,v)$ and an initial distribution $\nu$ for generating the first single state, we can get a Markov sequence $\{\tilde{\bm{u}}_{k}\}_{k\geq 1}$ generated by $\tilde{\bm{P}}_{2}$ with $\tilde{\bm{u}}_{1}\sim\tilde{\nu}_{2}$, where
$\tilde{\nu}_{2}\left(\widetilde{(u_{1},u_{2})}\right)=\begin{cases}\nu(u_{1})p(u_{1},u_{2})+\nu(u_{2})p(u_{1},u_{2}),\quad&\text{if}\,\ u_{1}\neq u_{2},\\ \nu(u_{1})p(u_{1},u_{2}),\quad&\text{if}\,\ u_{1}=u_{2}.\end{cases}$ (3.8)
3.1.3. Invariant distributions
The next lemma shows that $\mu_{2}$ and $\tilde{\mu}_{2}$ are indeed the invariant distributions of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$ respectively, and that $\tilde{\bm{P}}_{2}$ is a reversible transition matrix.
Lemma 3.1.
Let $\bm{P}$ be an irreducible transition matrix on $\Omega$ with an invariant distribution $\mu$. Then $\mu_{2}$ is an invariant distribution of $\bm{P}_{2}$, i.e.
$\mu_{2}(\bm{u})=\sum\limits_{\bm{v}\in\Omega^{2}}\mu_{2}(\bm{v})p_{2}(\bm{v},\bm{u}),\quad\text{for any}\,\ \bm{u}\in\Omega^{2}.$ (3.9)
Moreover, when $\bm{P}$ is reversible, $\tilde{\mu}_{2}$ is an invariant distribution of $\tilde{\bm{P}}_{2}$, and $\tilde{\bm{P}}_{2}$ satisfies the detailed balance condition:
$\tilde{\mu}_{2}(\tilde{\bm{u}})\tilde{p}_{2}(\tilde{\bm{u}},\tilde{\bm{v}})=\tilde{\mu}_{2}(\tilde{\bm{v}})\tilde{p}_{2}(\tilde{\bm{v}},\tilde{\bm{u}}),\quad\text{for any}\,\ \tilde{\bm{u}},\tilde{\bm{v}}\in\tilde{\Omega}^{2}.$ (3.10)
Proof.
Since $\mu$ is the invariant distribution of $\bm{P}$, we know (2.1) holds for any $u,v\in\Omega$. Thus, for any $\bm{u}=(u_{1},u_{2})\in\Omega^{2}$, we can compute that
$\sum\limits_{\bm{v}\in\Omega^{2}}\mu_{2}(\bm{v})p_{2}(\bm{v},\bm{u})$ $=\sum\limits_{v\in\Omega}\mu_{2}\left(v,u_{1}\right)p\left(u_{1},u_{2}\right)$ (3.11)
$=\sum\limits_{v\in\Omega}\mu(v)p\left(v,u_{1}\right)p\left(u_{1},u_{2}\right)$
$=\mu\left(u_{1}\right)p\left(u_{1},u_{2}\right)$
$=\mu_{2}(\bm{u}),$
which implies $\mu_{2}$ is the invariant distribution of $\bm{P}_{2}$.
When $\bm{P}$ is reversible, since (3.10) implies that $\tilde{\mu}_{2}$ is an invariant distribution of $\tilde{\bm{P}}_{2}$, we only need to show the reversibility of $\tilde{\bm{P}}_{2}$ under $\tilde{\mu}_{2}$. Notice that (3.10) always holds for $\tilde{\bm{u}}=\tilde{\bm{v}}$. On the other hand, for any $\tilde{\bm{u}}\neq\tilde{\bm{v}}$ which satisfy $\tilde{p}_{2}(\tilde{\bm{u}},\tilde{\bm{v}})\neq 0$, by the definition of $\tilde{\bm{P}}_{2}$, we can assume $\tilde{\bm{u}}\sim(u,w)$, $\tilde{\bm{v}}\sim(v,w)$, $u,v,w\in\Omega$. We then use the reversibility of $\bm{P}$ (2.4) to deduce that
$\tilde{\mu}_{2}(\tilde{\bm{u}})\tilde{p}_{2}(\tilde{\bm{u}},\tilde{\bm{v}})$ $=\mu(u)p(u,w)p(w,v)=\mu(v)p(v,w)p(w,u)=\tilde{\mu}_{2}(\tilde{\bm{v}})\tilde{p}_{2}(\tilde{\bm{v}},\tilde{\bm{u}}),$ (3.12)
which implies the reversibility of $\tilde{\bm{P}}_{2}$. ∎
3.2. Unbiased estimation
Now that we have formally introduced the induced Markov chains on path spaces, we can elaborate the connection between MLE and $\bm{P}_{2}$. By (1.1) and the definition of $\bm{P}_{2}$, the MLE estimator $\bm{R}_{n}=[r_{n}(u,v)]_{u,v\in\Omega}$ can be identified as a matrix empirical mean generated by $\bm{P}_{2}$:
$\bm{R}_{n}=\frac{1}{n}\sum\limits_{k=1}^{n}\bm{F}(\bm{u}_{k}),$ (3.13)
where $\bm{F}:\Omega^{2}\rightarrow\mathbb{R}^{d\times d}$ is a matrix-valued function defined by
$\bm{F}(\bm{w}):=\bm{E}_{w_{1}w_{2}},\quad\text{for any}\,\ \bm{w}=(w_{1},w_{2})\in\Omega^{2}.$ (3.14)
That is, the $(u,v)$ entry of $\bm{F}$ is just the indicator function $\mathbb{1}_{(u,v)}$ of the pair $(u,v)\in\Omega^{2}$, i.e. $\bm{F}=[\mathbb{1}_{(u,v)}]_{u,v\in\Omega}$.
Our SCE is associated with $\tilde{\bm{P}}_{2}$ in a similar way. More precisely, given a Markov sequence $\{\tilde{\bm{u}}_{k}\}_{k\geq 1}$ generated by $\tilde{\bm{P}}_{2}$, we can approximate $\bm{D}_{\mu}\bm{P}$ using the corresponding matrix empirical mean
$\bm{H}_{n}=\frac{1}{n}\sum\limits_{k=1}^{n}\tilde{\bm{F}}(\tilde{\bm{u}}_{k}),$ (3.15)
where $\tilde{\bm{F}}:\tilde{\Omega}^{2}\rightarrow\mathbb{H}_{d}$ is a matrix valued function defined by
$\tilde{\bm{F}}(\tilde{\bm{w}}):=\frac{1}{2}\left(\bm{E}_{w_{1}w_{2}}+\bm{E}_{w_{2}w_{1}}\right),\quad\text{for any}\,\ \tilde{\bm{w}}\sim(w_{1},w_{2}).$ (3.16)
Let $\tilde{f}_{uv}$ denote the $(u,v)$ entry of $\tilde{\bm{F}}$. Then
$\tilde{f}_{uv}=\tilde{f}_{vu}=\begin{cases}\frac{1}{2}\mathbb{1}_{\widetilde{(u,v)}},&\text{if}\,\ u\neq v,\\ \mathbb{1}_{\widetilde{(u,v)}},&\text{if}\,\ u=v,\end{cases}$ (3.17)
where $\mathbb{1}_{\widetilde{(u,v)}}$ is the indicator function of the symmetric pair $\widetilde{(u,v)}$.
The next lemma verifies that MLE and SCE are unbiased when the corresponding Markov chains on path spaces start from their invariant distributions.
Lemma 3.2.
It holds that
$\mathbb{E}_{\mu_{2}}\left[\bm{F}\right]=\bm{D}_{u}\bm{P}.$ (3.18)
Furthermore, when $\bm{P}$ is reversible,
$\mathbb{E}_{\tilde{\mu}_{2}}\left[\tilde{\bm{F}}\right]=\bm{D}_{u}\bm{P}.$ (3.19)
Proof.
We use the definition of $\bm{F}=[\mathbb{1}_{(u,v)}]_{u,v\in\Omega}$ to compute that
$\mathbb{E}_{\mu_{2}}\left[\mathbb{1}_{(u,v)}\right]=\mu_{2}(u,v)=\mu(u)p(u,v),$ (3.20)
which is the entry-wise form of (3.18).
When $\bm{P}$ satisfies the detailed balance condition (2.4), for any $u,v\in\Omega$ such that $u\neq v$, we have
$\operatorname{\mathbb{E}}_{\tilde{\mu}_{2}}\left[\tilde{f}_{uv}\right]=\frac{1}{2}\tilde{\mu}_{2}(\widetilde{(u,v)})=\frac{1}{2}(\mu(u)p(u,v)+\mu(v)p(v,u))=\mu(u)p(u,v).$ (3.21)
Moreover, for any $u\in\Omega$,
$\operatorname{\mathbb{E}}_{\tilde{\mu}_{2}}\left[\tilde{f}_{uu}\right]=\tilde{\mu}_{2}(\widetilde{(u,u)})=\mu(u)p(u,u).$ (3.22)
The above calculations imply (3.19). ∎
In the next section, we will study the convergence of $\bm{R}_{n}$ and $\bm{H}_{n}$ and establish non-asymptotic error bounds for MLE and SCE, before which we respectively propose the detailed frameworks of the two methods in Algorithm 1 and Algorithm 2 as follows.
Input: A sequence $\{u_{k}\}$ generated by $\bm{P}$ with length $N+1$;
1 $k=1$;
2 $\bm{R}=\sum\limits_{k=1}^{N}(\bm{E}_{u_{k}u_{k+1}})/N$;
3 $\hat{\mu}=\bm{R}\bm{1}$;
$\hat{\bm{P}}=\bm{D}_{\hat{\mu}}^{-1}\bm{R}$
Algorithm 1 Maximal Likelihood Estimation
Input: An oracle $\mathcal{P}$ of a reversible Markov chain, initial distribution $\nu$, total iteration number $N$;
1 $k=1$;
2 generate $u$ according to $\nu$;
3 $v=\mathcal{P}(u)$;
4 $\bm{H}=(\bm{E}_{uv}+\bm{E}_{vu})/(2N)$;
5 for $k=2:N$ do
6       generate $r\sim\mathcal{U}[0,1]$;
7       if $r\leq\frac{1}{2}$ then
8            $u\leftarrow\mathcal{P}(v)$;
9       end if
10      else
11            $v\leftarrow\mathcal{P}(u)$;
12       end if
13      $\bm{H}\leftarrow\bm{H}+(\bm{E}_{uv}+\bm{E}_{vu})/(2N)$;
14       $k\leftarrow k+1$;
15 end for
[…]
$\hat{\bm{P}}=\bm{D}_{\hat{\mu}}^{-1}\bm{H}$
Algorithm 2 Symmetric Counting Estimation
3.3. Spectral properties of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$
In order to derive non-asymptotic error bounds for MLE and SCE, we need to study the spectral gaps of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$, respectively. To start, the next theorem focuses on the spectral properties of $\bm{P}_{2}$. We will show that $\bm{P}$ and $\bm{P}_{2}$ share the same nonzero eigenvalues with algebraic multiplicity accounted by decomposing $\bm{P}_{2}$ as the product of two low-rank matrices according to the construction of $\bm{P}_{2}$. Furthermore, in order to obtain the spectral gaps of $\bm{P}_{2}$ required for concentration inequalities, we also study the spectrum of $\bm{P}_{2}\bm{P}_{2}^{*}$ and that of $\bm{P}_{2}+\bm{P}_{2}^{*}$. We find that, the absolute spectral gap of $\bm{P}_{2}$ is always zero, but the symmetric spectral gap of $\bm{P}_{2}$ is at least half of the symmetric spectral gap of $\bm{P}$, which makes it possible to apply certain concentration inequalities to $\bm{P}_{2}$ and derive non-asymptotic tail bounds for MLE.
Theorem 3.3.
Let $\bm{P}$ be an irreducible transition matrix on $\Omega$ with an invariant distribution $\mu$. The followings hold:
  1. (1)
    $\bm{P}$ and $\bm{P}_{2}$ share the same nonzero eigenvalues with algebraic multiplicity accounted.
  2. (2)
    $1$ is the only nonzero singular value of the matrix $\bm{D}_{\mu_{2}}^{1/2}\bm{P}_{2}\bm{D}_{\mu_{2}}^{-1/2}$ and the multiplicity of $1$ is $d$. As a result, $\eta_{a}(\bm{P}_{2})\equiv 0$.
  3. (3)
    It holds that
    $\eta_{s}(\bm{P}_{2})\geq\frac{\eta_{s}(\bm{P})}{2}>0,$ (3.23)
    and hence, $\mu_{2}$ is the unique invariant distribution of $\bm{P}_{2}$.
  4. (4)
    When $\bm{P}$ is reversible, any nonzero eigenvalue of $(\bm{P}_{2}+\bm{P}^{*}_{2})/2$ must be a nonzero eigenvalue of $(\bm{P}+\bm{\mathbf{I}})/2$ or $(\bm{P}-\bm{\mathbf{I}})/2$ with multiplicity accounted, and vise versa. As a result, $\eta_{s}(\bm{P}_{2})=\eta(\bm{P})/2$.
Proof.
From the construction of $\bm{P}_{2}$ we find that $\bm{P}_{2}$ can be decomposed as the product of two transition matrices that represent step 1 and step 2 introduced in Subsection 3.1.1, respectively. Let $\bm{S}$ denote the transition matrix of step 1 in which a given state $\bm{u}=(u_{1},u_{2})\in\Omega^{2}$ will jump to $u_{2}$ directly, i.e. for any $\bm{u}=(u_{1},u_{2})\in\Omega^{2}$, $v\in\Omega$,
$\bm{S}(\bm{u},v)=\delta_{u_{2}v},$ (3.24)
where $\delta_{uv}$ represents the Dirac symbol:
$\delta_{uv}=\begin{cases}1,\quad\text{if}\,\ u=v,\\ 0,\quad\text{if}\,\ u\neq v.\end{cases}$ (3.25)
Let $\bm{T}$ denote the transition matrix of step 2 in which a given state $u\in\Omega$ will jump to the state $(u,v)\in\Omega^{2}$ according to the probability
{choose $\mathbb{P}\left\{\text{choose }v\right\}=p(u,v),$ (3.26)
i.e. for any $u\in\Omega$, $\bm{v}=(v_{1},v_{2})\in\Omega^{2}$,
$\bm{T}(u,\bm{v})=p(u,v_{2})\delta_{uv_{1}}=p(v_{1},v_{2})\delta_{uv_{1}}.$ (3.27)
By the definition of $\bm{P}_{2}$ we know $\bm{P}_{2}=\bm{S}\bm{T}$. On the other hand, we find that, for any $u,v\in\Omega$,
$\bm{T}\bm{S}(u,v)=\sum\limits_{\bm{u}\in\Omega^{2}}\bm{T}(u,\bm{u})\bm{S}(\bm{u},v)=p(u,v),$ (3.28)
that is, $\bm{T}\bm{S}=\bm{P}$. We therefore deduce that $\bm{P}$ and $\bm{P}_{2}$ share the same nonzero eigenvalues from the basic fact that $\bm{S}\bm{T}$ and $\bm{T}\bm{S}$ share the same nonzero eigenvalues.
In order to obtain the absolute spectral gap of $\bm{P}_{2}$, we need to study the singular value of the matrix $\bm{D}_{\mu_{2}}^{1/2}\bm{P}_{2}\bm{D}_{\mu_{2}}^{-1/2}$. Let $\bm{S}_{1}=\bm{D}_{\mu_{2}}^{1/2}{\bm{S}}\bm{D}_{\mu}^{-1/2}$ and $\bm{T}_{1}=\bm{D}_{\mu}^{1/2}{\bm{T}}\bm{D}_{\mu_{2}}^{-1/2}$. From Theorem 3.3 we know
$\bm{S}_{1}\bm{T}_{1}=\bm{D}_{\mu_{2}}^{1/2}\bm{P}_{2}\bm{D}_{\mu_{2}}^{-1/2},\quad\bm{T}_{1}\bm{S}_{1}=\bm{D}_{\mu}^{1/2}\bm{P}\bm{D}_{\mu}^{-1/2}.$ (3.29)
For any $u,v\in\Omega$, we use the definition of the invariant distribution $\mu$ (2.1) to compute that
$\left({\bm{S}_{1}}^{\mathrm{T}}{\bm{S}_{1}}\right)(u,v)$ $=\mu(u)^{-1/2}\mu(v)^{-1/2}\sum\limits_{\bm{u}=(u_{1},u_{2})\in{\Omega}^{2}}{\bm{S}}(\bm{u},u){\mu}_{2}(\bm{u}){\bm{S}}(\bm{u},v)$ (3.30)
$=\mu(u)^{-1/2}\mu(v)^{-1/2}\sum\limits_{\bm{u}=(u_{1},u_{2})\in{\Omega}^{2}}\delta_{u_{2}u}\delta_{u_{2}v}{\mu}_{2}(\bm{u})$
$=\delta_{uv}\mu(u)^{-1}\sum\limits_{u_{1}\in{\Omega}}\mu(u_{1})p(u_{1},u)$
$=\delta_{uv},$
and
$\left({\bm{T}_{1}}{\bm{T}_{1}}^{\mathrm{T}}\right)(u,v)$ $=\mu(u)^{1/2}\mu(v)^{1/2}\sum\limits_{\bm{u}=(u_{1},u_{2})\in{\Omega}^{2}}{\bm{T}}(u,\bm{u}){\mu}_{2}(\bm{u})^{-1}{\bm{T}}(v,\bm{u})$ (3.31)
$=\mu(u)^{1/2}\mu(v)^{1/2}\sum\limits_{\bm{u}=(u_{1},u_{2})\in{\Omega}^{2}}\delta_{uu_{1}}\delta_{vu_{1}}p(u,u_{2})p(v,u_{2}){\mu}_{2}(\bm{u})^{-1}$
$=\delta_{uv}\mu(u)\sum\limits_{u_{2}\in\Omega}\frac{p(u,u_{2})^{2}}{\mu(u)p(u,u_{2})}$
$=\delta_{uv}.$
We thus deduce $\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}=\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}=\bm{\mathbf{I}}$. Since $\bm{S}_{1}\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}$ and $\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}=\bm{\mathbf{I}}$ share the same nonzero eigenvalues, we conclude that $1$ is the only nonzero singular value of $\bm{D}_{\mu_{2}}^{1/2}\bm{P}_{2}\bm{D}_{\mu_{2}}^{-1/2}$ and the multiplicity of $1$ is $d$. Thus $\lambda_{a}(\bm{P}_{2})=1$ and $\eta_{a}(\bm{P}_{2})=1-\lambda_{a}(\bm{P}_{2})=0$.
Next, we turn to study the symmetric spectral gap of $\bm{P}_{2}$. Recall that $\lambda_{s}(\bm{P}_{2})$ is the second largest eigenvalue of the matrix $(\bm{S}_{1}\bm{T}_{1}+\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}})/2=:\bm{K}$. Let us also define $\bm{L}:=(\bm{T}_{1}^{\mathrm{T}}\bm{T}_{1}+\bm{S}_{1}\bm{S}_{1}^{\mathrm{T}})/2$. To study the spectrum of $\bm{K}$, we first compute that
$\frac{1}{2}\left(\bm{K}+\bm{L}\right)=\frac{1}{4}\left(\bm{S}_{1}\bm{T}_{1}+\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\right)+\frac{1}{4}\left(\bm{T}_{1}^{\mathrm{T}}\bm{T}_{1}+\bm{S}_{1}\bm{S}_{1}^{\mathrm{T}}\right)=\frac{1}{4}\left(\bm{T}_{1}^{\mathrm{T}}+\bm{S}_{1}\right)\left(\bm{T}_{1}+\bm{S}_{1}^{\mathrm{T}}\right)\succeq 0,$ (3.32)
and
$\frac{1}{2}\left(\bm{K}-\bm{L}\right)=\frac{1}{4}\left(\bm{S}_{1}\bm{T}_{1}+\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\right)-\frac{1}{4}\left(\bm{T}_{1}^{\mathrm{T}}\bm{T}_{1}+\bm{S}_{1}\bm{S}_{1}^{\mathrm{T}}\right)=-\frac{1}{4}\left(\bm{T}_{1}^{\mathrm{T}}-\bm{S}_{1}\right)\left(\bm{T}_{1}-\bm{S}_{1}^{\mathrm{T}}\right)\preceq 0.$ (3.33)
As a result, we have
$\frac{1}{2}\left(\bm{K}-\bm{L}\right)\preceq\bm{K}=\frac{1}{2}\left(\bm{K}-\bm{L}\right)+\frac{1}{2}\left(\bm{K}+\bm{L}\right)\preceq\frac{1}{2}\left(\bm{K}+\bm{L}\right).$ (3.34)
which implies the eigenvalues of $\bm{K}$, in non-decreasing order, are less than or equal to the corresponding eigenvalues of $(\bm{K}+\bm{L})/2$ and larger than or equal to those of $(\bm{K}-\bm{L})/2$. On the other hand, by exchanging the order of multiplication, we have
$\frac{1}{4}\left(\bm{T}_{1}+\bm{S}_{1}^{\mathrm{T}}\right)\left(\bm{T}_{1}^{\mathrm{T}}+\bm{S}_{1}\right)$ $=\frac{1}{4}\left(\bm{T}_{1}\bm{S}_{1}+\bm{S}_{1}^{\mathrm{T}}\bm{T}_{1}^{\mathrm{T}}\right)+\frac{1}{4}\left(\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}+\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}\right)$ (3.35)
$=\frac{1}{2}\left(\bm{D}_{\mu}^{1/2}\frac{\bm{P}+\bm{\mathbf{I}}}{2}\bm{D}_{\mu}^{-1/2}+\bm{D}_{\mu}^{-1/2}\frac{\bm{P}^{\mathrm{T}}+\bm{\mathbf{I}}}{2}\bm{D}_{\mu}^{1/2}\right),$
and
$-\frac{1}{4}\left(\bm{T}_{1}-\bm{S}_{1}^{\mathrm{T}}\right)\left(\bm{T}_{1}^{\mathrm{T}}-\bm{S}_{1}\right)=\frac{1}{2}\left(\bm{D}_{\mu}^{1/2}\frac{\bm{P}-\bm{\mathbf{I}}}{2}\bm{D}_{\mu}^{-1/2}+\bm{D}_{\mu}^{-1/2}\frac{\bm{P}^{\mathrm{T}}-\bm{\mathbf{I}}}{2}\bm{D}_{\mu}^{1/2}\right).$ (3.36)
For any $\bm{M}\in\mathbb{H}_{d}$, we use $\lambda_{2}(\bm{M})$ to denote the second largest eigenvalue of $\bm{M}$. Since $(\bm{T}_{1}+\bm{S}_{1}^{\mathrm{T}})(\bm{T}_{1}^{\mathrm{T}}+\bm{S}_{1})/4$ and $(\bm{T}_{1}^{\mathrm{T}}+\bm{S}_{1})(\bm{T}_{1}+\bm{S}_{1}^{\mathrm{T}})/4$ share the same nonzero eigenvalues, we deduce that
$\lambda_{s}(\bm{P}_{2})=\lambda_{2}(\bm{K})\leq\lambda_{2}\left(\frac{\bm{K}+\bm{L}}{2}\right)=\lambda_{2}\left(\bm{D}_{\mu}^{1/2}\frac{\bm{P}+\bm{\mathbf{I}}}{4}\bm{D}_{\mu}^{-1/2}+\bm{D}_{\mu}^{-1/2}\frac{\bm{P}^{\mathrm{T}}+\bm{\mathbf{I}}}{4}\bm{D}_{\mu}^{1/2}\right)=\lambda_{s}\left(\frac{\bm{P}+\bm{\mathbf{I}}}{2}\right),$ (3.37)
and therefore
$\eta_{s}(\bm{P}_{2})=1-\lambda_{s}(\bm{P}_{2})\geq 1-\lambda_{s}\left(\frac{\bm{P}+\bm{\mathbf{I}}}{2}\right)=\frac{\eta_{s}(\bm{P})}{2}>0,$ (3.38)
where the last inequality is due to the irreducibility of $\bm{P}$. As a direct corollary, we know $1$ is a simple eigenvalue of $\bm{P}_{2}$ since
$\langle h,\left(\bm{\mathbf{I}}-\bm{P}_{2}\right)h\rangle_{\mu_{2}}\geq\eta_{s}(\bm{P}_{2})\left\|h\right\|_{\mu_{2}}>0,\qquad\text{for any }h\in L_{2,\mu_{2}}^{0}\backslash\{\bm{0}\},$ (3.39)
which implies that $\mu_{2}$ is the unique invariant distribution of $\bm{P}_{2}$.
Finally, when $\bm{P}$ is reversible, we have
$\bm{T}_{1}\bm{S}_{1}=\bm{D}_{\mu}^{1/2}\bm{P}\bm{D}_{\mu}^{-1/2}=\bm{D}_{\mu}^{-1/2}\bm{P}^{\mathrm{T}}\bm{D}_{\mu}^{1/2}=\bm{S}_{1}^{\mathrm{T}}\bm{T}_{1}^{\mathrm{T}},$ (3.40)
thus,
$\bm{K}^{2}$ $=\frac{1}{4}\left(\bm{S}_{1}\bm{T}_{1}\bm{S}_{1}\bm{T}_{1}+\bm{S}_{1}\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}+\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}\bm{T}_{1}+\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\right)$ (3.41)
$=\frac{1}{4}\left(\bm{S}_{1}\bm{T}_{1}\bm{S}_{1}\bm{T}_{1}+\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}+\bm{T}_{1}^{\mathrm{T}}\bm{T}_{1}+\bm{S}_{1}\bm{S}_{1}^{\mathrm{T}}\right)$
$=\frac{1}{4}\left(\bm{S}_{1}\bm{S}_{1}^{\mathrm{T}}\bm{T}_{1}^{\mathrm{T}}\bm{T}+\bm{T}_{1}^{\mathrm{T}}\bm{T}_{1}\bm{S}_{1}\bm{S}_{1}^{\mathrm{T}}+\bm{T}_{1}^{\mathrm{T}}\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}\bm{T}_{1}+\bm{S}_{1}\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}\bm{S}_{1}^{\mathrm{T}}\right)$
$=\bm{L}^{2}.$
By the fact that $\bm{K}\in\mathbb{H}_{d}$, $\bm{L}\in\mathbb{H}_{d}^{+}$, we know $\bm{L}$ must be the unique square root of $\bm{K}^{2}$, which implies that if $(\lambda,\alpha)$ is an eigen-pair of $\bm{K}$, then $(|\lambda|,\alpha)$ is an eigen-pair of $\bm{L}$. Therefore, any nonzero eigenvalue of $\bm{K}$ must be a nonzero eigenvalue of $(\bm{K}+\bm{L})/2$ or $(\bm{K}-\bm{L})/2$, and vice versa. Moreover, since $(\bm{K}-\bm{L})/2\preceq 0$ and $(\bm{K}-\bm{L})/2\succeq 0$, we know that if $\lambda_{2}((\bm{K}+\bm{L})/2)>0$, then it must also be the second largest eigenvalue of $\bm{K}$. On the other hand, if $\lambda_{2}((\bm{K}+\bm{L})/2)=0$, then 1 is the only nonzero eigenvalue of $(\bm{K}+\bm{L})/2$, and thus $\operatorname{rank}((\bm{K}+\bm{L})/2)=1$. We also know that
$\operatorname{rank}\left(\frac{\bm{K}-\bm{L}}{2}\right)=\operatorname{rank}\left(-\frac{1}{4}\left(\bm{T}_{1}^{\mathrm{T}}-\bm{S}_{1}\right)\left(\bm{T}_{1}-\bm{S}_{1}^{\mathrm{T}}\right)\right)\leq\operatorname{rank}\left(\bm{T}_{1}^{\mathrm{T}}-\bm{S}_{1}\right)\leq d,$ (3.42)
Therefore,
$\operatorname{rank}(\bm{K})=\operatorname{rank}\left(\frac{\bm{K}+\bm{L}}{2}+\frac{\bm{K}-\bm{L}}{2}\right)\leq\operatorname{rank}\left(\frac{\bm{K}+\bm{L}}{2}\right)+\operatorname{rank}\left(\frac{\bm{K}-\bm{L}}{2}\right)\leq 1+d<d^{2},$ (3.43)
which implies that 0 must be an eigenvalue of $\bm{K}$. To summarize, the following equality always holds:
$\lambda_{2}(\bm{K})=\lambda_{2}\left(\frac{\bm{K}+\bm{\mathbf{I}}}{2}\right).$ (3.44)
To complete the proof, notice that in the reversible case, (3.35) and (3.36) respectively reduce to
$\frac{1}{4}\left(\bm{T}_{1}+\bm{S}_{1}^{\mathrm{T}}\right)\left(\bm{T}_{1}^{\mathrm{T}}+\bm{S}_{1}\right)=\bm{D}_{\mu}^{1/2}\frac{\bm{P}+\bm{\mathbf{I}}}{2}\bm{D}_{\mu}^{-1/2},$ (3.45)
and
$\frac{1}{4}\left(\bm{T}_{1}-\bm{S}_{1}^{\mathrm{T}}\right)\left(\bm{T}_{1}^{\mathrm{T}}-\bm{S}_{1}\right)=\bm{D}_{\mu}^{1/2}\frac{\bm{P}-\bm{\mathbf{I}}}{2}\bm{D}_{\mu}^{-1/2}.$ (3.46)
As a result, any nonzero eigenvalue of $\bm{K}$ must be a nonzero eigenvalue of $(\bm{P}+\bm{\mathbf{I}})/2$ or $(\bm{P}-\bm{\mathbf{I}})/2$ with multiplicity accounted, and vise versa. Besides, by (3.45), we have
$\lambda_{s}(\bm{P}_{2})=\lambda_{2}(\bm{K})=\lambda_{2}\left(\frac{\bm{K}+\bm{L}}{2}\right)=\lambda\left(\frac{\bm{P}+\bm{\mathbf{I}}}{2}\right),$ (3.47)
which implies $\eta_{s}(\bm{P}_{2})=\eta(\bm{P})/2$. This completes the proof. ∎
Next, we study the spectrum of $\tilde{\bm{P}}_{2}$ when $\bm{P}$ is reversible. Compared with Theorem 3.3, we can obtain the spectral gaps of $\tilde{\bm{P}}_{2}$ through a simper approach since $\tilde{\bm{P}}_{2}$ is also a reversible transition matrix. The next theorem shows that $\tilde{\bm{P}}_{2}$ and $(\bm{P}+\bm{\mathbf{I}})/2$ share the same nonzero eigenvalues with multiplicity accounted. As a result, the spectral gaps of $\tilde{\bm{P}}_{2}$ can be obtained directly.
Theorem 3.4.
Let $\bm{P}$ be an irreducible reversible transition matrix on $\Omega$ with an invariant distribution $\mu$. Then $(\bm{P}+\bm{\mathbf{I}})/2$ and $\tilde{\bm{P}}_{2}$ share the same nonzero eigenvalues with multiplicity accounted. As a result, $\tilde{\mu}_{2}$ is the unique invariant distribution of $\tilde{\bm{P}}_{2}$. Moreover, it holds that
$\eta_{a}\left(\tilde{\bm{P}}_{2}\right)=\eta\left(\tilde{\bm{P}}_{2}\right)=\frac{\eta(\bm{P})}{2}>0.$ (3.48)
Proof.
Similar to the proof for $\bm{P}_{2}$, we decompose $\tilde{\bm{P}}_{2}$ as the product of two transition matrices which represent step 1 and step 2 introduced in Subsection 3.1.2, respectively. To be specific, let $\tilde{\bm{S}}$ denote the transition matrix of step 1 in which a given state $\tilde{\bm{u}}\sim(u,v)$ will jump to $u$ or $v$ with the equal probability $1/2$, i.e. for any $\tilde{\bm{u}}\in\tilde{\Omega}^{2}$,$u\in\Omega$,
$\tilde{\bm{S}}(\tilde{\bm{u}},u)=\begin{cases}1/2,&\tilde{\bm{u}}\sim(u,v),\text{where }v\neq u,\\ 1,&\tilde{\bm{u}}\sim(u,u),\\ 0,&\text{otherwise}.\end{cases}$ (3.49)
Let $\tilde{\bm{T}}$ denote the transition matrix of step 2 in which a given state $u$ will jump to the state $\tilde{\bm{v}}\sim(u,v)$ according to the probability
{choose $\mathbb{P}\left\{\text{choose }v\right\}=p(u,v),$ (3.50)
i.e. for any $u\in\Omega$, $\tilde{\bm{v}}\in\tilde{\Omega}^{2}$,
$\tilde{\bm{T}}(u,\tilde{\bm{v}})=\begin{cases}p(u,v),&\tilde{\bm{v}}\sim(u,v),v\in\Omega,\\ 0,&\text{otherwise}.\end{cases}$ (3.51)
Then we have $\tilde{\bm{P}}_{2}=\tilde{\bm{S}}\tilde{\bm{T}}$. Similarly, for any $u,v\in\Omega$ such that $u\neq v$, we have
$\tilde{\bm{T}}\tilde{\bm{S}}(u,v)=\sum\limits_{\tilde{\bm{u}}\in\tilde{\Omega}^{2}}\tilde{\bm{T}}(u,\tilde{\bm{u}})\tilde{\bm{S}}(\tilde{\bm{u}},v)=\frac{1}{2}p(u,v).$ (3.52)
On the other hand, direct calculation shows that
$\tilde{\bm{T}}\tilde{\bm{S}}(u,u)$ $=\sum\limits_{\tilde{\bm{u}}\in\tilde{\Omega}^{2}}\tilde{\bm{T}}(u,\tilde{\bm{u}})\tilde{\bm{S}}(\tilde{\bm{u}},u)=\sum\limits_{\tilde{\bm{u}}\sim(u,v),v\in\Omega}\tilde{\bm{T}}(u,\tilde{\bm{u}})\tilde{\bm{S}}(\tilde{\bm{u}},u)$ (3.53)
$=\sum\limits_{v\neq u}\frac{1}{2}p(u,v)+p(u,u)=1+\frac{1}{2}p(u,u).$
Thus $\tilde{\bm{T}}\tilde{\bm{S}}=(\bm{P}+\bm{\mathbf{I}})/2$, which implies that $(\bm{P}+\bm{\mathbf{I}})/2$ and $\tilde{\bm{P}}_{2}$ share the same nonzero eigenvalues with multiplicity accounted. As a direct corollary, by the irreducibility of $\bm{P}$ we know 1 is a simple eigenvalue of $\bm{P}$, and therefore 1 is also a simple eigenvalue of $\tilde{\bm{P}}_{2}$, which establishes the uniqueness of the invariant distribution of $\tilde{\bm{P}}_{2}$.
Furthermore, by the reversibility of $\tilde{\bm{P}}_{2}$ and the fact that all eigenvalues of $\tilde{\bm{P}}_{2}$ are nonnegative, we deduce
$\lambda_{a}\left(\tilde{\bm{P}}_{2}\right)=\lambda\left(\tilde{\bm{P}}_{2}\right)=\lambda\left(\frac{\bm{P}+\bm{\mathbf{I}}}{2}\right)=\frac{1+\lambda(\bm{P})}{2},$ (3.54)
which implies
$\eta_{a}\left(\tilde{\bm{P}}_{2}\right)=\eta\left(\tilde{\bm{P}}_{2}\right)=1-\frac{1+\lambda(\bm{P})}{2}=\frac{\eta(\bm{P})}{2}.$ (3.55)
This completes the proof. ∎
4. Non-asymptotic error bounds
With the spectral properties of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$ in hand, we can now prove the main results of this paper by establishing non-asymptotic error bounds for MLE and SCE, respectively.
4.1. Tail bounds of MLE and SCE
We will first derive tail bounds for the error of MLE and SCE using concentration inequalities for Markov chains. It is remarkable that most of the existing results about concentration inequalities for Markov chains typically require a nonzero absolute spectral gap of the transition matrix, which are inapplicable to $\bm{P}_{2}$ since $\eta_{a}(\bm{P}_{2})=0$. Therefore, we choose to adopt the following Bernstein-type inequality obtained in [8], which only requires a nonzero symmetric spectral gap of the transition matrix.
Theorem 4.1.
Let $\{u_{k}\}_{k\geq 1}$ be a Markov chain on $\Omega$ generated by a transition kernel $\bm{P}$ with an invariant distribution $\mu$ and a symmetric spectral gap $\eta_{s}(\bm{P})>0$, and let $g:\Omega\rightarrow\mathbb{R}$ be a scalar function that satisfies $\mathbb{E}_{\mu}[g]=0$, $\mathbb{E}_{\mu}[g^{2}]\leq\sigma^{2}$, and $|g|\leq M$. It holds that for any $t>0$,
$\mathbb{P}\left\{\left|\frac{1}{n}\sum\limits_{j=1}^{n}g(u_{j})\right|\geq t\right\}\leq 2\|\nu/\mu\|_{\infty}\exp\left(\frac{-n\eta_{s}(\bm{P})t^{2}}{4M\sqrt{\left(2+5\eta_{s}(\bm{P})/\sqrt{2}\right)^{2}\sigma^{2}+t^{2}}}\right),$ (4.1)
where $\nu$ is the initial distribution of $u_{1}$.
We then use Theorem 4.1 to derive non-asymptotic tail bounds for MLE.
Proof of Theroem 2.1 (1)(2).
Recall that in MLE, the corresponding matrix empirical mean $\bm{R}_{n}$ can be reformulated as
$\bm{R}_{n}=\frac{1}{n}\sum\limits_{k=1}^{n}\bm{F}(\bm{u}_{k}),$ (4.2)
where $\{\bm{u}_{k}\}_{k\geq 1}$ is the Markov chain constructed in Subsection 3.1.1, and $\bm{F}=[\mathbb{1}_{(u,v)}]_{u,v\in\Omega}$ is the matrix-valued function whose $(u,v)$ entry is the indicator function $\mathbb{1}_{(u,v)}$ of the pair $(u,v)\in\Omega^{2}$. It is shown that
$\operatorname{\mathbb{E}}_{\mu_{2}}\left[\bm{F}\right]=\bm{D}_{\mu}\bm{P}.$ (4.3)
in Lemma 3.2, and thus the almost surely convergence of $\bm{R}_{n}$ is a direct corollary of the standard ergodic theorem for Markov chains.
Next, for any $u,v\in\Omega$, we take $g=\mathbb{1}_{(u,v)}-\mu(u)p(u,v)$. Then it is easy to verify that $\mathbb{E}_{\mu}[g]=0$, $|g|\leq 1$, and $\mathbb{E}_{\mu}[g^{2}]\leq 1$. We therefore substitute $M=\sigma^{2}=1$, $\eta_{s}(\bm{P}_{2})\geq\eta_{s}(\bm{P})/2$ into Theorem 4.1 and deduce
$\mathbb{P}\left\{\left|r_{n}(u,v)-\mu(u)p(u,v)\right|\geq t\right\}\leq 2\|\nu_{2}/\mu_{2}\|_{\infty}\exp\left(\frac{-n\eta_{s}(\bm{P})t^{2}}{8\sqrt{\left(2+5\eta_{s}(\bm{P})/2\sqrt{2}\right)^{2}+t^{2}}}\right),$ (4.4)
where $\nu_{2}$ is the distribution of $\bm{u}_{1}$. Finally, for any $\bm{w}=(w_{1},w_{2})\in\Omega^{2}$, we have
$\frac{\nu_{2}(\bm{w})}{\mu_{2}(\bm{w})}=\frac{\nu(w_{1})p(w_{1},w_{2})}{\mu(w_{1})p(w_{1},w_{2})}=\frac{\nu(w_{1})}{\mu(w_{1})},$ (4.5)
which implies $\|\nu_{2}/\mu_{2}\|_{\infty}=\|\nu/\mu\|_{\infty}$. The proof is completed. ∎
As for SCE, we choose to derive non-asymptotic tail bounds directly in the matrix form using matrix concentration inequalities. Recently, [16] established Hoeffding and Bernstein-type matrix concentration inequalities (in $l_{2}$ operator norm) for sums of Markov dependent random matrices, under the assumption that the transition matrix admits a nonzero absolute spectral gap. We present a simplified version of [16, Theorem 2.2] as follows.
Theorem 4.2.
Let $\{u_{k}\}_{k\geq 1}$ be a Markov chain on $\Omega$ generated by a transition kernel $\bm{P}$ with an invariant distribution $\mu$ and an absolute spectral gap $\eta_{a}(\bm{P})>0$, and let $\bm{G}:\Omega\rightarrow\mathbb{H}_{d}$ be a matrix-valued function that satisfies $\mathbb{E}_{\mu}[\bm{G}]=0$, $\|\mathbb{E}_{\mu}[\bm{G}^{2}]\|\leq\sigma^{2}$, and $\|\bm{G}\|\leq M$. It holds that for any $t>0$,
$\mathbb{P}\left\{\left\|\frac{1}{n}\sum\limits_{j=1}^{n}\bm{G}(u_{j})\right\|\geq t\right\}\leq 2\|\nu/\mu\|_{\infty}d^{2-\frac{\pi}{4}}\exp\left(\frac{-n\eta_{a}(\bm{P})t^{2}}{32\pi^{-2}(2-\eta_{a}(\bm{P}))\sigma^{2}+256\pi^{-3}Mt}\right),$ (4.6)
where $\nu$ is the initial distribution of $u_{1}$.
Since we have shown that $\eta_{a}(\tilde{\bm{P}}_{2})=\eta_{a}(\bm{P})/2$, we can use Theorem 4.2 to establish non-asymptotic tail bounds for SCE.
Proof of Theroem 2.2 (1)(2).
In SCE, the corresponding empirical mean $\bm{H}_{n}$ is defined by
$\bm{H}_{n}=\frac{1}{n}\sum\limits_{k=1}^{n}\tilde{\bm{F}}(\tilde{\bm{u}}_{k}),$ (4.7)
where $\{\tilde{\bm{u}}_{k}\}_{k\geq 1}$ is a Markov chain generated by $\tilde{\bm{P}}_{2}$, and $\tilde{\bm{F}}:\tilde{\Omega}^{2}\rightarrow\mathbb{H}_{d}$ is the matrix-valued function on the symmetric path space $\tilde{\Omega}^{2}$ which satisfies
$\tilde{\bm{F}}(\tilde{\bm{w}})=\frac{1}{2}\left(\bm{E}_{w_{1}w_{2}}+\bm{E}_{w_{2}w_{1}}\right),\qquad\text{ if}\,\ \tilde{\bm{w}}\sim(w_{1},w_{2}).$ (4.8)
It is shown that
$\operatorname{\mathbb{E}}_{\tilde{\mu}_{2}}\left[\tilde{\bm{F}}\right]=\bm{D}_{\mu}\bm{P}$ (4.9)
in Lemma 3.2, and then the almost surely convergence of $\bm{H}_{n}$ follows immediately. Next, we apply Theorem 4.2 with $\bm{G}=\tilde{\bm{F}}-\bm{D}_{\mu}\bm{P}$. We find that
$\left\|\bm{G}\right\|\leq\left\|\tilde{\bm{F}}\right\|+\left\|\bm{D}_{\mu}\bm{P}\right\|\leq 2,$ (4.10)
and
$\left\|\mathbb{E}_{\tilde{\mu}_{2}}\left[\bm{G}^{2}\right]\right\|=\left\|\mathbb{E}_{\tilde{\mu}_{2}}\left[\tilde{\bm{F}}^{2}\right]-\left(\bm{D}_{\mu}\bm{P}\right)^{2}\right\|\leq\left\|\mathbb{E}_{\tilde{\mu}_{2}}\left[\tilde{\bm{F}}^{2}\right]\right\|+\left\|\bm{D}_{\mu}\bm{P}\right\|^{2}\leq 2.$ (4.11)
Substituting $M=\sigma^{2}=2$, $\eta_{a}(\bm{\tilde{P}_{2}})=\eta(\bm{P})/2$ into Theorem 4.2, we have
$\mathbb{P}\left\{\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|\geq t\right\}\leq 2\|\tilde{\nu}_{2}/\tilde{\mu}_{2}\|_{\infty}d^{2-\frac{\pi}{4}}\exp\left(\frac{-n\eta(\bm{P})t^{2}}{128\pi^{-2}(2-\eta(\bm{P})/2)+1024\pi^{-3}t}\right),$ (4.12)
where $\tilde{\nu}_{2}$ is the distribution of $\tilde{\bm{u}}_{1}$. According to the generation of $\tilde{\bm{u}}_{1}$, for any $\tilde{\bm{w}}\sim(w_{1},w_{2})$, we can compute that
$\frac{\tilde{\nu}_{2}\left(\tilde{\bm{w}}\right)}{\tilde{\mu}_{2}\left(\tilde{\bm{w}}\right)}=\frac{\nu(w_{1})p(w_{1},w_{2})+\nu(w_{2})p(w_{2},w_{1})}{\mu(w_{1})p(w_{1},w_{2})+\mu(w_{2})p(w_{2},w_{1})}=\frac{1}{2}\left(\frac{\nu(w_{1})}{\mu(w_{1})}+\frac{\nu(w_{2})}{\mu(w_{2})}\right).$ (4.13)
Therefore, we have $\|\tilde{\nu}_{2}/\tilde{\mu}_{2}\|_{\infty}=\|\nu/\mu\|_{\infty}$, which implies the claimed tail bound. ∎
Remark 4.3.
The reason we can establish matrix-form tail bounds for SCE is that $\tilde{\bm{P}}_{2}$ always admits a nonzero absolute spectral gap as long as $\bm{P}$ is irreducible. In contrast, the absolute spectral gap of $\bm{P}_{2}$ is always zero, which makes the result proposed in [16] inapplicable to MLE.
4.2. Dimension-free expectation bound
In this subsection, we derive dimension-free expectation bounds for the errors of MLE and SCE and complete the proofs of Theorem 2.1 and Theorem 2.2.
Proof of Theorem 2.1 (3).
First, let $\nu$ be the initial distribution of $u_{1}$, and let $\nu_{2}$ be the initial distribution induced by $\nu$. Notice that
$\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|u_{1}\sim\nu\right]$ $=\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\bm{u}_{1}\sim\nu_{2}\right]$ (4.14)
$\leq\|\nu_{2}/\mu_{2}\|_{\infty}\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\bm{u}_{1}\sim\mu_{2}\right]$
$=\|\nu/\mu\|_{\infty}\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\bm{u}_{1}\sim\mu_{2}\right].$
Therefore, we only need to bound $\mathbb{E}[\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\|_{\mathrm{F}}^{2}]$ under the condition that $\bm{u}_{1}\sim\mu_{2}$. For any $\bm{v}\in\Omega^{2}$, let $\bar{\mathbb{1}}_{\bm{v}}$ denote the projection of the indicator function $\mathbb{1}_{\bm{v}}$ onto the mean-zero subspaces $L_{2,\mu_{2}}^{0}$, i.e.
$\bar{\mathbb{1}}_{\bm{v}}(\bm{w})=\mathbb{1}_{\bm{v}}(\bm{w})-\mu_{2}(\bm{v}),\quad\text{for any}\,\ \bm{w}\in\Omega^{2}.$ (4.15)
we then use the definition of $\bm{R}_{n}$ to compute that
$\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]$ $=\mathbb{E}\left[\sum\limits_{\bm{v}\in\Omega^{2}}\left(\frac{1}{n}\sum\limits_{k=1}^{n}\bar{\mathbb{1}}_{\bm{v}}\left(\bm{u}_{k}\right)\right)^{2}\middle|\bm{u}_{1}\sim\mu_{2}\right]$
$=\frac{1}{n^{2}}\sum\limits_{\bm{v}\in\Omega^{2}}\sum\limits_{1\leq i,j\leq n}\mathbb{E}\left[\bar{\mathbb{1}}_{\bm{v}}\left(\bm{u}_{i}\right)\cdot\bar{\mathbb{1}}_{\bm{v}}\left(\bm{u}_{j}\right)\middle|\bm{u}_{1}\sim\mu_{2}\right]$
$=\frac{1}{n^{2}}\sum\limits_{\bm{v}\in\Omega^{2}}\left(n\left\|\bar{\mathbb{1}}_{\bm{v}}\right\|^{2}_{\mu_{2}}+2\sum\limits_{1\leq i<j\leq n}\mathbb{E}\left[\bar{\mathbb{1}}_{\bm{v}}\left(\bm{u}_{i}\right)\cdot\bm{P}_{2}^{j-i}\bar{\mathbb{1}}_{\bm{v}}\left(\bm{u}_{j}\right)\middle|\bm{u}_{i}\sim\mu_{2}\right]\right)$
$=\frac{1}{n^{2}}\sum\limits_{\bm{v}\in\Omega^{2}}\left(n\left\|\bar{\mathbb{1}}_{\bm{v}}\right\|^{2}_{\mu_{2}}+2\sum\limits_{k=1}^{n-1}(n-k)\left\langle\bar{\mathbb{1}}_{\bm{v}},\bm{P}_{2}^{k}\bar{\mathbb{1}}_{\bm{v}}\right\rangle_{\mu_{2}}\right)$
$=\frac{1}{n}\left(\sum\limits_{\bm{v}\in\Omega^{2}}\left\|\bar{\mathbb{1}}_{\bm{v}}\right\|^{2}_{\mu_{2}}\right)+\frac{2}{n^{2}}\sum\limits_{\bm{v}\in\Omega^{2}}\sum\limits_{k=1}^{n-1}(n-k)\left\langle\bar{\mathbb{1}}_{\bm{v}},\left(\bm{P}_{2}^{k}-\bm{\Pi}_{2}\right)\bar{\mathbb{1}}_{\bm{v}}\right\rangle_{\mu_{2}}$
$\leq\left(\frac{1}{n}+\frac{2}{n^{2}}\left\|\sum\limits_{k=1}^{n-1}(n-k)\left(\bm{P}_{2}^{k}-\bm{\Pi}_{2}\right)\right\|_{\mu_{2}}\right)\left(\sum\limits_{\bm{v}\in\Omega^{2}}\left\|\bar{\mathbb{1}}_{\bm{v}}\right\|^{2}_{\mu_{2}}\right),$
where $\bm{\Pi}_{2}:=\bm{1}_{d^{2}}\mu_{2}^{\mathrm{T}}$ is the projection operator onto $\bm{1}_{d^{2}}$ under the $L_{2,\mu_{2}}$ inner product. Since
$\sum\limits_{\bm{v}\in\Omega^{2}}\left\|\bar{\mathbb{1}}_{\bm{v}}\right\|^{2}_{\mu_{2}}\leq\sum\limits_{\bm{v}\in\Omega^{2}}\left\|\mathbb{1}_{\bm{v}}\right\|^{2}_{\mu_{2}}=\sum\limits_{\bm{v}\in\Omega^{2}}\mu_{2}(\bm{v})=1,$ (4.16)
we only need to bound the $L_{2,\mu_{2}}$ operator norm of $\sum_{k=1}^{n-1}(n-k)(\bm{P}_{2}^{k}-\bm{\Pi}_{2})$. Let $\bm{S}$, $\bm{T}$ be defined in Theorem 3.3 which satisfy $\bm{P}_{2}=\bm{S}\bm{T}$, $\bm{P}=\bm{T}\bm{S}$, and let $\bm{S}_{1}=\bm{D}_{\mu_{2}}^{1/2}{\bm{S}}\bm{D}_{\mu}^{-1/2}$, $\bm{T}_{1}=\bm{D}_{\mu}^{1/2}{\bm{T}}\bm{D}_{\mu_{2}}^{-1/2}$. One can verify that $\bm{S}\bm{1}_{d}=\bm{1}_{d^{2}}$ and $\mu^{\mathrm{T}}\bm{T}=\mu_{2}^{\mathrm{T}}$, and thus
$\bm{\Pi}_{2}=\bm{1}_{d^{2}}\mu_{2}^{\mathrm{T}}=\bm{S}\bm{1}_{d}\mu^{\mathrm{T}}\bm{T}=\bm{S}\bm{\Pi}\bm{T}.$ (4.17)
where $\bm{\Pi}$ is the projection onto $\bm{1}_{d}$ defined in Subsection 2.1. By (2.6), we know that $\bm{P}^{k}-\bm{\Pi}=(\bm{P}-\bm{\Pi})^{k}=(\bm{P}-\bm{\Pi})^{k}(\bm{\mathbf{I}}-\bm{\Pi})$ holds for any $k\in\mathbb{N}_{+}$. As a result, for any integer $k\geq 2$, we have
$\bm{P}_{2}^{k}-\bm{\Pi}_{2}=\left(\bm{S}\bm{T}\right)^{k}-\bm{S}\bm{\Pi}\bm{T}=\bm{S}\left(\left(\bm{T}\bm{S}\right)^{k-1}-\bm{\Pi}\right)\bm{T}=\bm{S}\left(\bm{P}-\bm{\Pi}\right)^{k-1}(\bm{\mathbf{I}}-\bm{\Pi})\bm{T}.$ (4.18)
In particular, $\bm{P}_{2}-\bm{\Pi}_{2}=\bm{S}(\bm{\mathbf{I}}-\bm{\Pi})\bm{T}$. For convenience, we define a polynomial
$\varphi(x):=\frac{2}{n^{2}}\sum\limits_{k=1}^{n-1}(n-k)x^{k-1}.$ (4.19)
Then
$\frac{2}{n^{2}}\left\|\sum\limits_{k=1}^{n-1}(n-k)\left(\bm{P}_{2}^{k}-\bm{\Pi}_{2}\right)\right\|_{\mu_{2}}$ $=\frac{2}{n^{2}}\left\|\bm{S}\left((n-1)\bm{\mathbf{I}}+\sum\limits_{k=2}^{n-1}(n-k)\left(\bm{P}-\bm{\Pi}\right)^{k-1}\right)(\bm{\mathbf{I}}-\bm{\Pi})\bm{T}\right\|_{\mu_{2}}$ (4.20)
$=\left\|\bm{S}\varphi\left(\bm{P}-\bm{\Pi}\right)(\bm{\mathbf{I}}-\bm{\Pi})\bm{T}\right\|_{\mu_{2}}$
$=\left\|\bm{S}_{1}\bm{D}_{\mu}^{1/2}\varphi\left(\bm{P}-\bm{\Pi}\right)(\bm{\mathbf{I}}-\bm{\Pi})\bm{D}_{\mu}^{-1/2}\bm{T}_{1}\right\|$
$\leq\|\bm{S}_{1}\|\|\bm{T}_{1}\|\left\|\bm{D}_{\mu}^{1/2}\varphi\left(\bm{P}-\bm{\Pi}\right)(\bm{\mathbf{I}}-\bm{\Pi})\bm{D}_{\mu}^{-1/2}\right\|$
$=\left\|\varphi\left(\bm{P}-\bm{\Pi}\right)(\bm{\mathbf{I}}-\bm{\Pi})\right\|_{\mu},$
where the last equality is due to the fact that $\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}=\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}=\bm{\mathbf{I}}$ (Theorem 3.3). Our goal is to bound $\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})\|_{\mu}$. To achieve this, consider an arbitrary $h\in L_{2,\mu}$, and let $\bar{h}:=(\bm{\mathbf{I}}-\bm{\Pi})h$. If $\bar{h}=\bm{0}$, then $\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})h=\bm{0}$. Otherwise,
$\frac{\left\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})h\right\|_{\mu}}{\left\|h\right\|_{\mu}}\leq\frac{\left\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})h\right\|_{\mu}}{\left\|\bar{h}\right\|_{\mu}}=\frac{\left\|\varphi(\bm{P}-\bm{\Pi})\bar{h}\right\|_{\mu}}{\left\|\bar{h}\right\|_{\mu}}=\frac{\left\|\varphi(\bm{P})\bar{h}\right\|_{\mu}}{\left\|\bar{h}\right\|_{\mu}}.$ (4.21)
As a result,
$\left\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})\right\|_{\mu}$ $=\sup\limits_{h\in L_{2,\mu},\ h\neq\bm{0}}\frac{\left\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})h\right\|_{\mu}}{\left\|h\right\|_{\mu}}=\sup\limits_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\left\|\varphi(\bm{P})h\right\|_{\mu}}{\left\|h\right\|_{\mu}}.$ (4.22)
By the irreducibility of $\bm{P}$, we have $\operatorname{rank}(\bm{\mathbf{I}}-\bm{P})=d-1$. We also know that for any $h\in L_{2,\mu}$, $\operatorname{\mathbb{E}}_{\mu}[(\bm{\mathbf{I}}-\bm{P})h]=0$, and that $(\bm{\mathbf{I}}-\bm{P})\bm{1}=0$. Thus, $\bm{\mathbf{I}}-\bm{P}$ is a bijection from $L_{2,\mu}^{0}$ to $L_{2,\mu}^{0}$. Hence, for any $h\in L_{2,\mu}^{0}$, there exists a function $g\in L_{2,\mu}^{0}$ such that $(\bm{\mathbf{I}}-\bm{P})g=h$. For such $g$, by the definition (2.8) of the symmetric spectral gap $\eta_{s}(\bm{P})$ , we have
$\|g\|_{\mu}^{2}\leq\frac{\left\langle g,(\bm{\mathbf{I}}-\bm{P})g\right\rangle_{\mu}}{\eta_{s}(\bm{P})}=\frac{\left\langle g,h\right\rangle_{\mu}}{\eta_{s}(\bm{P})}\leq\frac{\|g\|_{\mu}\|h\|_{\mu}}{\eta_{s}(\bm{P})},$ (4.23)
which implies $\|g\|_{\mu}\leq\|h\|_{\mu}/\eta_{s}(\bm{P})$. Therefore,
$\left\|\varphi(\bm{P})h\right\|_{\mu}=\left\|\varphi(\bm{P})(\bm{\mathbf{I}}-\bm{P})g\right\|_{\mu}\leq\left\|\varphi(\bm{P})(\bm{\mathbf{I}}-\bm{P})\right\|_{\mu}\left\|g\right\|_{\mu}\leq\frac{\left\|\varphi(\bm{P})(\bm{\mathbf{I}}-\bm{P})\right\|_{\mu}}{\eta_{s}(\bm{P})}\|h\|_{\mu}.$ (4.24)
To continue, notice that
$\left\|\varphi(\bm{P})(\bm{\mathbf{I}}-\bm{P})\right\|_{\mu}=\frac{2}{n^{2}}\left\|(n-1)\bm{\mathbf{I}}-\sum\limits_{k=1}^{n-1}\bm{P}^{k}\right\|_{\mu}\leq\frac{2}{n^{2}}\left((n-1)+\sum\limits_{k=1}^{n-1}\left\|\bm{P}^{k}\right\|_{\mu}\right)\leq\frac{4}{n}.$ (4.25)
We thus deduce
$\left\|\varphi(\bm{P})h\right\|_{\mu}\leq\frac{4}{n\eta_{s}(\bm{P})}\left\|h\right\|_{\mu},\quad\text{for any}\,\ h\in L_{2,\mu}^{0},$ (4.26)
which implies $\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})\|_{\mu}\leq 4/(n\eta_{s}(\bm{P})$. As a result,
$\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\leq\left(\frac{1}{n}+\left\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})\right\|_{\mu}\right)\left(\sum\limits_{\bm{v}\in\Omega^{2}}\left\|\bar{\mathbb{1}}_{\bm{v}}\right\|^{2}_{\mu_{2}}\right)\leq\frac{1}{n}+\frac{4}{n\eta_{s}(\bm{P})}.$ (4.27)
which is the claimed dimension-free expectation bound.
If $\bm{P}$ admits a nonzero absolute spectral gap $\eta_{a}(\bm{P})=1-\lambda_{a}(\bm{P})$, we can bound $\|\varphi(\bm{P}-\bm{\Pi})\|_{\mu}$ alternatively using the fact that
$\left|\varphi(x)\right|=\frac{2}{n^{2}}\left|\sum\limits_{k=1}^{n-1}(n-k)x^{k-1}\right|=\frac{2}{n^{2}}\left(\frac{n}{1-x}+\frac{x^{n}-1}{(1-x)^{2}}\right)\leq\frac{2}{n(1-x)},\quad\text{for any}\,\ x\in(-1,1).$ (4.28)
Since all coefficients of $\varphi$ are nonnegative, we have
$\left\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})\right\|_{\mu}\leq\left\|\varphi\left(\bm{P}-\bm{\Pi}\right)\right\|_{\mu}$ $\leq\varphi\left(\left\|\bm{P}-\bm{\Pi}\right\|_{\mu}\right)=\varphi\left(\lambda_{a}(\bm{P})\right)\leq\frac{2}{n\eta_{a}(\bm{P})}.$ (4.29)
It follows that,
$\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\leq\frac{2+\eta_{a}(\bm{P})}{n\eta_{a}(\bm{P})}.$ (4.30)
Moreover, when $\bm{P}$ is reversible, $\bm{P}-\bm{\Pi}$ is self-adjoint under the $L_{2,\mu}$ inner product, and $\|\varphi(\bm{P}-\bm{\Pi})\|_{\mu}$ is the largest singular value of the symmetric matrix $\varphi(\bm{D}_{\mu}^{1/2}(\bm{P}-\bm{\Pi})\bm{D}_{\mu}^{-1/2})$. In this case, (4.29) can be improved to
$\left\|\varphi\left(\bm{P}-\bm{\Pi}\right)\right\|_{\mu}$ $=\left\|\varphi\left(\bm{D}_{\mu}^{1/2}(\bm{P}-\bm{\Pi})\bm{D}_{\mu}^{-1/2}\right)\right\|$ (4.31)
$=\sup\limits_{\lambda\in\operatorname{spec}\left(\bm{P}-\bm{\Pi}\right)}|\varphi(\lambda)|\leq\sup\limits_{\lambda\in\operatorname{spec}\left(\bm{P}-\bm{\Pi}\right)}\frac{2}{n(1-\lambda)}=\frac{2}{n\eta(\bm{P})},$
which implies
$\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\leq\frac{2+\eta(\bm{P})}{n\eta(\bm{P})}.$ (4.32)
This completes the proof of Theorem 2.1. ∎
Remark 4.4.
Compared with (4.27), the error bound in (4.30) has a better constant. Therefore, (4.30) could be sharper than (4.27) when $\eta_{a}(\bm{P})\geq\eta_{s}(\bm{P})/2$. However, we cannot guarantee that $\bm{P}$ admits a nonzero absolute spectral gap $\eta_{a}(\bm{P})$. In contrast, $\eta_{s}(\bm{P})>0$ holds for any irreducible transition matrix $\bm{P}$. From this point of view, (4.27) is a more robust result.
Remark 4.5.
(4.27) can be view as a matrix version of [4, Theorem 1.2]. In fact, one can apply [4, Theorem 1.2] to $\{\bm{u}_{k}\}_{k\geq 1}$ to obtain an error bound that involves the symmetric spectral gap of $\bm{P_{2}}$. As for our results, we utilize the special structure of $\bm{P}_{2}$ in the proof so that we can use the spectral properties of $\bm{P}$ directly to obtain a sharper error bound.
Proof of Theroem 2.2 (3).
Let $\tilde{\nu}_{2}$ be the initial distribution induced by $\nu$. Since
$\mathbb{E}\left[\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\tilde{\bm{u}}_{1}\sim{\tilde{\nu}_{2}}\right]$ $\leq\|\tilde{\nu}_{2}/\tilde{\mu}_{2}\|_{\infty}\mathbb{E}\left[\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\tilde{\bm{u}}_{1}\sim{\tilde{\mu}_{2}}\right]$ (4.33)
$=\|{\nu}/{\mu}\|_{\infty}\mathbb{E}\left[\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\tilde{\bm{u}}_{1}\sim{\tilde{\mu}_{2}}\right],$
it suffices to bound $\mathbb{E}[\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\|_{\mathrm{F}}^{2}]$ under the condition that $\tilde{\bm{u}}_{1}\sim{\tilde{\mu}_{2}}$. For any $\tilde{\bm{v}}\in\Omega^{2}$, we use $\bar{\mathbb{1}}_{\tilde{\bm{v}}}$ to denote the projection of the indicator function $\mathbb{1}_{\tilde{\bm{v}}}$ onto the mean-zero subspaces $L_{2,\tilde{\mu}_{2}}^{0}$, i.e.
$\bar{\mathbb{1}}_{\tilde{\bm{v}}}({\tilde{\bm{w}}})=\mathbb{1}_{\tilde{\bm{v}}}(\tilde{\bm{w}})-\tilde{\mu}_{2}(\tilde{\bm{v}}),\quad\text{for any}\,\ \tilde{\bm{w}}\in\Omega^{2}.$ (4.34)
By the definition of $\bm{H}_{n}=[h_{n}(u,v)]_{u,v\in\Omega}$, we have
$h_{n}(u,v)-\mu(u)p(u,v)=\begin{cases}\sum_{k=1}^{n}\bar{\mathbb{1}}_{\widetilde{(u,v)}}(\tilde{\bm{u}}_{k})/(2n)\quad&\text{if }\,\ u\neq v,\\ \sum_{k=1}^{n}\bar{\mathbb{1}}_{\widetilde{(u,v)}}(\tilde{\bm{u}}_{k})/n\quad&\text{if }\,\ u=v.\end{cases}$ (4.35)
Next, we define a polynomial $\tilde{\varphi}$ as
$\tilde{\varphi}(x):=\frac{1}{n}+\frac{2}{n^{2}}\sum\limits_{k=1}^{n-1}(n-k)x^{k},$ (4.36)
and then compute that
$\mathbb{E}\left[\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]$ $=\mathbb{E}\left[\sum\limits_{\begin{subarray}{c}u,v\in\Omega\\ u\neq v\end{subarray}}\left(\frac{1}{2n}\sum\limits_{k=1}^{n}\left({\bar{\mathbb{1}}_{\widetilde{(u,v)}}\left(\tilde{\bm{u}}_{k}\right)}\right)\right)^{2}+\sum\limits_{\begin{subarray}{c}u\in\Omega\end{subarray}}\left(\frac{1}{n}\sum\limits_{k=1}^{n}\left({\bar{\mathbb{1}}_{\widetilde{(u,u)}}\left(\tilde{\bm{u}}_{k}\right)}\right)\right)^{2}\middle|\tilde{\bm{u}}_{1}\sim{\tilde{\mu}_{2}}\right]$
$\leq\mathbb{E}\left[\sum\limits_{\tilde{\bm{v}}\in\tilde{\Omega}^{2}}\left(\frac{1}{n}\sum\limits_{k=1}^{n}\left({\bar{\mathbb{1}}_{\tilde{\bm{v}}}\left(\tilde{\bm{u}}_{k}\right)}\right)\right)^{2}\middle|\tilde{\bm{u}}_{1}\sim{\tilde{\mu}_{2}}\right]$
$=\frac{1}{n^{2}}\sum\limits_{\tilde{\bm{v}}\in\tilde{\Omega}^{2}}\sum\limits_{1\leq i,j\leq n}\mathbb{E}\left[\bar{\mathbb{1}}_{\tilde{\bm{v}}}\left(\tilde{\bm{u}}_{i}\right)\cdot\bar{\mathbb{1}}_{\tilde{\bm{v}}}\left(\tilde{\bm{u}}_{j}\right)\middle|\tilde{\bm{u}}_{1}\sim{\tilde{\mu}_{2}}\right]$
$=\sum\limits_{\tilde{\bm{v}}\in\tilde{\Omega}^{2}}\left(\frac{1}{n}\left\langle\bar{\mathbb{1}}_{\tilde{\bm{v}}},\bar{\mathbb{1}}_{\tilde{\bm{v}}}\right\rangle_{\tilde{\mu}_{2}}+\frac{2}{n^{2}}\sum\limits_{k=1}^{n-1}(n-k)\left\langle\left(\tilde{\bm{P}}_{2}^{k}-\tilde{\bm{\Pi}}_{2}\right)\bar{\mathbb{1}}_{\tilde{\bm{v}}},\bar{\mathbb{1}}_{\tilde{\bm{v}}}\right\rangle_{\tilde{\mu}_{2}}\right)$
$\leq\left\|\frac{\bm{\mathbf{I}}}{n}+\frac{2}{n^{2}}\sum\limits_{k=1}^{n-1}(n-k)\left(\tilde{\bm{P}}_{2}^{k}-\tilde{\bm{\Pi}}_{2}\right)\right\|_{\tilde{\mu}_{2}}\left(\sum\limits_{\tilde{\bm{v}}\in\tilde{\Omega}^{2}}\left\|\bar{\mathbb{1}}_{\tilde{\bm{v}}}\right\|_{\tilde{\mu}_{2}}^{2}\right)$
$\leq\left\|\tilde{\varphi}\left(\tilde{\bm{P}}_{2}-\tilde{\bm{\Pi}}_{2}\right)\right\|_{\tilde{\mu}_{2}},$
where $\tilde{\bm{\Pi}}_{2}$ is the projection operator onto $\bm{1}_{d(d+1)/2}$ under the $L_{2,\tilde{\mu}_{2}}$ inner product. For the last inequality, we use the facts that $\tilde{\bm{P}}_{2}^{k}-\tilde{\bm{\Pi}}_{2}=(\tilde{\bm{P}}_{2}-\tilde{\bm{\Pi}}_{2})^{k}$ and $\sum_{\tilde{\bm{v}}\in\tilde{\Omega}^{2}}\|\bar{\mathbb{1}}_{\tilde{\bm{v}}}\|_{\tilde{\mu}_{2}}^{2}\leq 1$. Notice that for any $x\in[0,1)$,
$\tilde{\varphi}(x)=\frac{1}{n}+\frac{2}{n^{2}}\sum\limits_{k=1}^{n-1}(n-k)x^{k}=\frac{1}{n^{2}}\left(\frac{n(1+x)}{1-x}+\frac{2x^{n+1}-2x}{\left(1-x\right)^{2}}\right)\leq\frac{1+x}{n(1-x)}.$ (4.37)
Therefore,
$\left\|\tilde{\varphi}\left(\tilde{\bm{P}}_{2}-\tilde{\bm{\Pi}}_{2}\right)\right\|_{\tilde{\mu}_{2}}$ $\leq\tilde{\varphi}\left(\left\|\tilde{\bm{P}}_{2}-\tilde{\bm{\Pi}}_{2}\right\|_{\tilde{\mu}_{2}}\right)=\tilde{\varphi}\left(\lambda_{a}\left(\tilde{\bm{P}}_{2}\right)\right)$ (4.38)
$\leq\frac{1+\lambda_{a}\left(\tilde{\bm{P}}_{2}\right)}{n\left(1-\lambda_{a}\left(\tilde{\bm{P}}_{2}\right)\right)}=\frac{2-\eta_{a}\left(\tilde{\bm{P}}_{2}\right)}{n\eta_{a}\left(\tilde{\bm{P}}_{2}\right)}=\frac{4-\eta(\bm{P})}{n\eta\left(\bm{P}\right)},$
where the last equality is due to the fact that $\eta_{a}(\tilde{\bm{P}}_{2})=\eta(\bm{P})/2$ (Theorem 3.4). As a result, we obtain
$\mathbb{E}\left[\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\leq\frac{4-\eta(\bm{P})}{n\eta\left(\bm{P}\right)},$ (4.39)
which completes the proof. ∎
Remark 4.6.
We can also use the tail bounds proved in Subsection 4.1 to derive expectation bounds for SCE. In general, Let $X$ be a random variable such that
$\mathbb{P}\left\{|X|\geq t\right\}\leq c_{1}\mathrm{e}^{-c_{2}t^{2}},\quad\text{for any}\,\ t>0,$ (4.40)
where $c_{1}\geq 1,c_{2}>0$ are two constants. Then $\mathbb{E}[X^{2}]$ can be bounded by integrating the tail bound:
$\mathbb{E}\left[X^{2}\right]=\int_{0}^{+\infty}\mathbb{P}\left\{|X|^{2}>t\right\}\mathrm{d}{t}\leq\int_{0}^{+\infty}\min\left\{1,c_{1}\mathrm{e}^{-c_{2}t}\right\}\mathrm{d}{t}=\frac{\log\left(c_{1}\mathrm{e}\right)}{c_{2}}.$ (4.41)
Thus we can deduce $\operatorname{\mathbb{E}}[\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\|^{2}]=\mathcal{O}(\log(d)/(n\eta_{a}(\bm{P})))$ from the tail bounds of SCE. However, from the proofs given in this subsection, we know the error expectations of the two methods are both independent of the dimension $d$.
5. numerical results
In this section, we present a numerical study to verify and visualize our theoretical results. In order to compare MLE and SCE, we take $\bm{P}$ to be a reversible transition matrix. Besides, we adopt the following strategy to adjust the spectral gap of $\bm{P}$: for $10<\eta<$ 1 and a reversible transition matrix $\bm{P}$ with an invariant distribution $\mu$, we take
$\bm{P}_{\eta}=\begin{cases}\frac{\eta}{1-\lambda_{2}}\bm{P}+\left(1-\frac{\eta}{1-\lambda_{2}}\right)\bm{\mathbf{I}},&\eta\leq 1-\lambda_{2},\\ \frac{\lambda_{2}}{1-\eta}\bm{P}+\left(1-\frac{\lambda_{2}}{1-\eta}\right)\bm{1}\mu^{\mathrm{T}},&\eta>1-\lambda_{2},\end{cases}$ (5.1)
where $\lambda_{2}$ is the second largest eigenvalue of $\bm{P}$. Then it is not hard to see that $\bm{P}_{\eta}$ is a reversible transition matrix with a spectral gap $\eta$. For a given transition matrix $\bm{P}$, we run MLE and SCE $N$ times to obtain $\|\bm{H}_{n}^{(k)}-\bm{D}_{\mu}\bm{P}\|_{\mathrm{F}}^{2}$ and $\|\bm{R}_{n}^{(k)}-\bm{D}_{\mu}\bm{P}\|_{\mathrm{F}}^{2}$, $k=1,2,\dots N$, and then average over $k$ to get the mean square error (MSE). Figure 5.1 and Figure 5.2 respectively show the linear dependence of 1/MSE on $\eta$ and $n$, while Figure 5.3 demonstrates that the MSEs of two methods are both independent of the size of the state space. We can also find out that $\text{MSE}_{\text{MLE}}/\text{MSE}_{\text{SCE}}\approx 1/2$ when the spectral gap of $\bm{P}$ is small, which supports the theoretical results presented in Theorem 2.1 and Theorem 2.2.
(a) $\eta(\bm{P})=0.5$
(b) $\eta(\bm{P})=0.2$
(c) $\eta(\bm{P})=0.01$
Figure 5.1. Plots of $1/\text{MSE}$ as the function of $n$ for $\eta(\bm{P})=0.5,0.2,0.01$, respectively. We fix $N=10000$ and $d=50$.
(a) $n=5000$
(b) $n=20000$
(c) $n=50000$
Figure 5.2. Plots of $1/\text{MSE}$ as the function of $\eta(\bm{P})$ for $n=5000,20000,50000$, respectively. We fix $N=10000$ and $d=50$.
Figure 5.3. Plot of MSE as the function of $d$ for $N=5000$, $n=1000$. For each $d\in\left\{10,20,\cdots 200\right\}$, we take the average of the MSEs of 100 different $d\times d$ transition matrices with fixed spectral gap $\eta=0.1$.
Acknowledgements.
The authors are supported by the National Key R&D Program of China under the grant 2021YFA1001500.